### v2

# DAG-Plan: Generating Directed Acyclic Dependency Graphs for Dual-Arm Cooperative Planning

Abstract
Dual-arm robots offer enhanced versatility and efficiency over single-arm counterparts by enabling concurrent manipulation of multiple objects or cooperative execution of tasks using both arms. However, effectively coordinating the two arms for complex long-horizon tasks remains a significant challenge. Existing task planning methods predominantly focus on single-arm robots or rely on predefined bimanual operations, failing to fully leverage the capabilities of dual-arm systems. To address this limitation, we introduce DAG-Plan, a structured task planning framework tailored for dual-arm robots. DAG-Plan harnesses large language models (LLMs) to decompose intricate tasks into actionable sub-tasks represented as nodes within a directed acyclic graph (DAG). Critically, DAG-Plan dynamically assigns these sub-tasks to the appropriate arm based on real-time environmental observations, enabling parallel and adaptive execution. We evaluate DAG-Plan on the novel Dual-Arm Kitchen Benchmark, comprising 9 sequential tasks with 78 sub-tasks and 26 objects. Extensive experiments demonstrate the superiority of DAG-Plan over directly using LLM to generate plans, achieving nearly 50% higher efficiency compared to the single-arm task planning baseline and nearly double the success rate of the dual-arm task planning baseline.
Keywords: Dual-arm Robots, Task Planning, Large Language Models
1 Introduction
Achieving effective bimanual coordination in robotics is challenging due to the complexities of dual-arm operations, requiring precise spatial and temporal coordination [1, 2]. While humans effortlessly coordinate their hands in daily tasks, replicating such coordination in robots presents significant challenges for both planning and learning-based methods. Traditionally, planning-based methods have focused on motion planning, employing hand-designed primitives to manage the movements of two robotic arms [3, 4]. However, these methods often fall short in dynamic or intricate environments as they lack the flexibility required for adaptive task execution. In contrast, learning-based strategies, such as Reinforcement Learning (RL) and Imitation Learning (IL), provide more adaptive solutions by enabling robots to learn control policies from either human-designed rewards [5, 6] or human demonstrations [7, 8]. However, these methods struggle to generalize to zero-shot scenarios, where robots need to execute tasks without prior specific training.
Large language models (LLMs) have emerged as powerful tools endowed with extensive knowledge and sophisticated reasoning abilities [9, 10, 11, 12]. By systematically breaking down tasks into actionable sub-tasks and leveraging their commonsense knowledge and implicit reasoning capabilities, LLMs empower robots to effectively adapt to new scenarios and tasks [13, 14, 15, 16, 17, 18]. However, existing planning methods are primarily applied to single-arm robots, focusing on using one arm to perform skills. While some studies [19, 20] have employed dual-arm robots as test platforms, these still engage only one arm at a time or rely on predefined bimanual tasks, leading to inefficiencies. In this work, we propose the development of dual-arm cooperative planning using LLMs. This approach facilitates parallel task execution, thereby significantly improving efficiency. Directly generating dual-arm planning schemes as subgoal sequences with LLMs faces significant challenges, primarily two gaps: 1) Dual-arm collaboration allows multiple sub-tasks to be executed simultaneously, making the temporal dependencies between them highly complex. Previous approaches, which employed linear temporal dependency lists for task sequencing, have proven to be inefficient for planning and execution; 2) The non-interactivity with the environment, as the execution order of the task sequence and the side of the executing arm are fixed, making it impossible to choose executable and cost-effective sub-tasks based on the environment and robot state during execution.
To address these challenges, we introduce DAG-Plan, a structured task planning framework that leverages the capabilities of LLMs. We leverage Directed Acyclic Graph (DAG) as a task graph for dual-arm task planning. A DAG represents each complex task as actionable sub-tasks, with nodes indicating these sub-tasks and directed edges defining explicit temporal dependencies. Firstly, we utilize LLMs to generate the DAG, decomposing complex tasks into nodes, each associated with a specific type and the number of arms required for execution. Subsequently, the DAG enters the task planning inference process. DAG-Plan uses this temporal dependency information and node types to determine priority candidate nodes and common candidate nodes, assigning them to the left and right arms. DAG-Plan checks the feasibility and calculates the cost of combinations of left and right arm candidate nodes based on the environmental state, adaptively executing sub-tasks that are easier and closer to perform.
Our main contributions are summarized as follows: 1) We present DAG-Plan, an efficient cooperative task planning framework for mobile dual-arm robots. This framework represents decomposed sub-tasks as a DAG and dynamically assigns these sub-tasks to the appropriate arm based on the real-time environment state; 2) We conduct the Dual-arm Kitchen Benchmark based on Sapien engine [21], which consists of 9 long-horizon tasks with 78 sub-tasks and 26 assets and objects. This Benchmark is manually constructed by robotic experts and includes plan tests and physical simulation tests; 3) Extensive experiments on the Dual-arm Kitchen Benchmark show that DAG-Plan significantly outperforms other methods. DAG-Plan achieves nearly a 50% increase in efficiency over the baseline by directly employing LLMs to generate the single-arm plan. Compared to the baseline, which directly utilizes LLMs for dual-arm plans, DAG-Plan demonstrates nearly double the success rate.
Figure 1: An overview of DAG-Plan. The DAG-Plan generates a DAG based on human instruction and environmental description. It checks the graph’s completeness and reflects the LLM to regenerate if incomplete. Once a valid DAG is obtained, DAG-Plan performs task inference to identify executable candidate nodes. The occupied arm and free arm are assigned priority candidate nodes and common candidate nodes respectively. The framework then evaluates all candidate combinations for feasibility and cost. DAG-Plan selects the nodes with the lowest cost and employs motion planning and reinforcement learning for execution. DAG-Plan updates the graph, iterating inference until the DAG is fully executed.
2 Related Works
Task Planning with LLMs. LLMs are increasingly being used to generate sequences of executable actions that enable an agent to achieve goals represented in natural language. Previous studies have successfully utilized the commonsense and in-context learning capabilities of pre-trained LLMs to create executable plans for embodied agents [22, 23, 19, 24, 25, 20, 26, 27, 28]. However, most of these studies have been applied to single-arm robots, only requiring consideration of executing a single-arm action at one timestep. Although LLM+P [19] and CoPAL [20] both have designed the tasks for dual-arm robots. The plan generated by LLM+P and CoPAL still can not manipulate two different objects with each arm at the same time causing low execution efficiency. Additionally, these studies [19, 27, 28] explicitly use the operational rules of the environment as input, whereas we only provide an environment description, relying on the LLM’s inherent world modeling capability.
Dual-arm Robot Manipulation. Dual-arm coordination has made progress in industrial [1, 2] and agricultural scenarios [29, 30] with fixed process operations and domestic settings [31, 32] with single-skill operations. With the development of methods such as motion planning [30, 32], reinforcement learning [33, 34], and imitation learning [35, 36], dual-arm robots can perform many human-like operations at the skill level. However, they still lack the ability to autonomously plan in zero-shot complex scenarios. The commonsense and contextual learning capabilities of pre-trained LLMs make it possible for dual-arm robots to autonomously plan in zero-shot complex scenarios.
Structured Task Decomposition (STD). STD involves breaking down a complex task into a DAG. Previous methods for STD, such as Crowd-Sourced STD [37, 38] and Query-based STD [39, 40], were limited by data availability. However, LLMs contain extensive real-world commonsense knowledge, offering new approaches for STD. TaskLAMA [41] has conducted detailed research on structured task decomposition using LLMs, demonstrating that LLMs can decompose real-world tasks into task graphs with temporal dependencies. In this work, we explore the use of DAG to address issues in dual-arm robot sequential planning with low execution efficiency.
3 Method
We utilize LLMs to generate a DAG, where each task for a dual-arm robot is represented as a node. The directed edges between these nodes are crucial as they establish a clear and mandatory sequence of tasks, dictating the order in which tasks must be performed. This ensures that dependencies are meticulously adhered to, allowing for efficient task execution. The robot dynamically assesses the state of its environment and the status of its arms to select the next optimal tasks from the graph. This ongoing selection process prioritizes activities, keeping both arms continuously engaged, either cooperatively or independently, depending on task requirements. By emphasizing the directed nature of these task sequences, the methodology enhances operational efficiency, minimizes idle time, and ensures a smooth workflow, significantly improving the robot’s performance in complex operational environments. An overview of the DAG-Plan pipeline is illustrated in Figure 1.
3.1 Directed Acyclic Sub-task Dependency Graph Generation
In prior research, task planning for robots typically involves generating a linear sequence of sub-tasks. However, this model falls short in scenarios involving dual-arm robots, where the capability for parallel task execution can significantly enhance operational efficiency. By coordinating both arms, many tasks can be performed concurrently, which the linear model does not exploit fully.
To address this limitation, we propose a novel approach using LLMs to generate optimized dual-arm plans. Our method involves decomposing complex tasks into a dual-arm task graph instead of a linear sequence. This graph better represents the complex temporal dependencies of bimanual operations. We define the graph as $G=(V,E,T,N)$, where $V$ denotes the tasks, $E$ represents the dependencies, $T$ categorizes the task types, and $N$ specifies the arm requirements. Each vertex $v_{i}\in V$ corresponds to a specific sub-task, and each directed edge $e_{ij}=(v_{i},v_{j})\in E$ indicates that $v_{i}$ must be completed before $v_{j}$. The task types include: 1) Occupy, where tasks involve the engagement of the robot’s gripper and the arm will be occupied after execution, typically for grasping or holding an object; 2) Tool use, which refers to tasks that require the use of a tool, remaining in the gripper throughout the operation; 3) Release, for tasks where an object is released from the gripper, often associated with placement or release into a specific location; 4) Operate, denoting general operational tasks that leave the gripper free post-completion, where the arm will be occupied during execution and released afterward; and 5) Complete, which marks the end of all tasks, represented as the terminal node in the graph. This classification aids in specifying the nature of the task and the number of arms required, thereby enabling a more sophisticated and efficient planning strategy tailored for dual-arm robotic systems. The arm number of node $n_{i}\in N$ represents the arm number of node needed. If $n_{i}=1$ a node requires one arm. If $n_{i}=2$, both arms are needed for execution. The occupy-release pairs are crucial structures in dual-arm task graphs. An occupy-release pair mainly consists of an occupy node as start point and a release node as end point, with potentially several tool use nodes in between. A complete occupy-release pair ensures that the robot arm is not continuously occupied and prevents placing an object without first grasping it. Additionally, the dual-arm task graph should be a fully connected DAG. After generating the dual-arm task graph, we check it for completeness. If the graph contains incomplete grasp-release pairs or is not fully connected, we reflect the LLMs to regenerate the task graph.
Figure 2: The process of Task Planning Inference. In the task “cut the carrot and red bell pepper", DAG-Plan initializes common candidate nodes based on the DAG. It evaluates node combinations, checks feasibility, and calculates costs. The right arm is selected to grasp the carrot and the left to grasp the red bell pepper. After execution, the task graph and nodes are updated, adding subsequent release nodes to the priority candidate nodes for each arm. In stage 2, each arms are assigned corresponding priority candidate nodes, checked, and executed. The task graph and nodes are updated again. The priority nodes become empty, indicating the arms are free. In stage 3, with only a single node left, the closest arm to the target executes this node.
3.2 Task Planning Inference with Generated Directed Acyclic Graph
After generating the dual-arm task graph, the planning process enters the inference phase. This phase utilizes the task graph and the observed state to dynamically refine the planning of robotic arm operations. It selects and executes sub-tasks that are executable and have the lowest cost, based on the task graph and the current environment state. As shown in Figure 2, we provide a detailed and specific illustration of the task planning inference process. We first identify the two types of candidate nodes in DAG-Plan: common candidate nodes and priority candidate nodes. The common candidate nodes include those that can be executed when the robotic arm is in the free state. The priority candidate nodes include subsequent nodes when the arm is already engaged in an occupy-release pair. The common candidate nodes are initially selected by identifying nodes within the graph that no other nodes point to.
During execution, once a node completes its operation, it and its associated edges are removed from the graph. This unlocks nodes that are dependent solely on the executed node. If the node executed involves operations like operate or release, the corresponding arm becomes free, and any dependent nodes unlocked by this action are added to the common candidates. Conversely, if the executed node involves actions like occupy or tool use, where the arm remains occupied, any dependent nodes in this occupy-release pair are unlocked and placed into the priority candidates for that arm. When the priority candidate nodes for a specific arm are not empty, the arm must select a sub-task from these priority candidate nodes. This ensures that the arm’s next tasks are aimed at completing actions necessary to free up the arm. When there are no more priority candidates for an arm, that arm is considered free and can select tasks from the common candidates. This strategic selection and execution framework ensures efficient operation and task handling by the dual-arm robotic system.
Once we obtain the candidate nodes for each robotic arm, we generate all possible combinations of left and right arm candidate nodes. These combinations are then checked for feasibility, and any pairs that fail the checks are removed from the candidate set. There are three checks in total. The first check involves verifying the presence of an occupy node within the candidates. If an occupy node is found, we further examine whether its successor nodes contain dependencies that require other conditions to be met first. This could lead to prolonged arm occupancy, thereby decreasing operational efficiency. The second check assesses the distance between target objects for the left and right arms. If these targets are beyond a specified distance threshold, it becomes difficult for both arms to operate simultaneously. The third check evaluates the relative positions of the target locations for each arm. If the target position for the left arm is to the right of the right arm’s target, there is a risk of the arms crossing and colliding, which would prevent the execution of these candidate nodes. After completing the checks, we calculate the cost of the left and right hand candidate nodes based on the environment state. We aim for the target objects to be close to the robotic arms and to each other, facilitating dual-arm operations. Therefore, the cost is composed of two parts: the distance from each target to the respective arm, and the distance between the targets. The pair with the lowest combined cost is selected for execution. The cost $J$ is represented as:
$J=\text{dis}(obj_{\text{right}},obj_{\text{left}})+\frac{1}{2}\left[\text{dis}(obj_{\text{right}},hand_{\text{right}})+\text{dis}(obj_{\text{left}},hand_{\text{left}})\right].$
3.3 Motion Planning and Reinforcement Learning Mixed Skill Learning
Once the sub-tasks to be executed are determined, the robot needs to perform the corresponding actions to bridge the gap between textual instructions and the physical environment. Reinforcement Learning (RL) is a powerful technique that enables robots to acquire and refine skills autonomously by interacting with their environment[42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. By engaging in a trial-and-error process and receiving feedback in the form of rewards, robots can learn to optimize their actions to achieve specific objectives [55, 56, 57, 58, 59, 60]. Recent advancements [61, 62, 63, 64] have further enhanced automation by using LLMs to generate text-to-reward mappings. We combine motion planning (RRT-connect) [65] and RL (PPO) [47] to facilitate skill learning (see details in Appendix A). The robot approaches the target object using motion planning and then learns the corresponding skill through RL. We input the generated dual-arm task graph and description of the reward APIs into LLMs, which then combine different reward APIs and determine if they are necessary conditions for completing the task, thus achieving skill learning. However, given the sensitivity of RL training and the ambiguity of language, RL policies may fail to achieve the goal or achieve it in unexpected ways. Therefore, we manually fine-tune the generated rewards to ensure that the final skills meet the task requirements.
4 Experiments
4.1 Experimental Setup
To validate the correctness and execution efficiency of our method, we created a Dual-arm Kitchen Benchmark, including plan tests and physical simulation tests. This benchmark fills the gap in long-sequence operation rigid body physics simulation benchmarks for dual-arm robots. The goal of this benchmark is to validate the success rate and efficiency of dual-arm robot planning in complex scenarios. The dual-arm robot should correctly plan and fully utilize both the left and right arms, completing tasks with as few execution stages as possible. The benchmark consists of 9 sequential tasks (see details in Appendix B) are shown in Table 1, comprising a total of 78 sub-tasks. The main setting is a modern kitchen divided into three scenarios shown in Figure 3, containing a total of 26 assets and objects. Our physical simulation scene is built on the Sapien [21]. The embodied platform for planning execution is a dual-arm robot with a mobile base, each 7-degree-of-freedom arm equipped with a two-finger gripper. The platform is equipped with five RGB-D cameras, attached above and below each arm’s gripper and on the robot’s head.
Table 1: Task list of Dual-arm Kitchen Benchmark.
Index Instruction Index Instruction
Task 1 Put the apple and bread onto the plate Task 6 Heat the soup and pour a cup of cola
Task 2 Place the apple on the plate and toast the bread Task 7 Put the apple and the pear into the bowl
Task 3 Juice the apple and toast the bread Task 8 Cut the carrot and red bell pepper
Task 4 Wash the cup and the bowl Task 9 Make a pot of soup
Task 5 Heat the soup and put the tin on the table
Figure 3: Snapshots of 3 Scenarios of Dual-arm Kitchen Benchmark. Scenario 1 encompasses tasks 1-3, scenario 2 encompasses tasks 4-6, and scenario 3 encompasses tasks 7-9.
Evaluation of Planning Effectiveness and Conciseness. In this experiment, we focus on testing the conciseness of the generated plans and the number of stages required. In a stage, the robot can execute a right arm node and a left arm node. This requires that the plans generated by the LLMs can achieve the task goals in terms of language logic and do not violate the preconditions for stage execution. Since the lower-level execution method uses reinforcement learning, learning from incorrect plans is costly and unproductive, making plan validation crucial. We used GPT-4111https://openai.com/api. This work mainly uses gpt-4-turbo-2024-04-09. under the temperature of 0.5 to generate five plans for each task, evaluating their Success Rate (SR) and the fewest number of Stage of the passed plan required at the language level. Finally, we calculated the average success rate and the average number of stages for all tasks. We defined Stage Efficiency as the ratio of single-arm plan stages to the stages required by each method. For failed dual-arm plans, we calculated the stage count based on the single-arm plan stages to ensure a fair comparison.
Evaluation with Physical Simulation. In this experiment, we will test the executability and execution efficiency of the plans in physical simulation scenarios. Compared to plan tests, physical simulation tests validate both the high-level planning and low-level execution capabilities. We will use reinforcement learning to acquire executable skills from the plans that passed the plan tests and then execute them in the physical simulation environment. We select the plan that passes the plan test and requires the minimum stages for skill learning. Then evaluate the Success Rate (SR) of execution and minimum execution Time in the physical environment with 10 trials. Finally, we calculated the average success rate and the execution time for all tasks. We defined Execution Efficiency as the ratio of single-arm plan execution time to the execution time required by each method. For failed dual-arm plans, we calculated the average time based on the single-arm execution time to ensure a fair comparison.
Task Planning for Single-arm (TP-S) directly uses LLMs to generate a full task list, with each stage involving a single arm or both arms to manipulate a single object. Task Planning for Dual-arm (TP-D) also directly uses LLMs to generate a full task list, but each stage can use the arms to manipulate either a single object or two different objects. Our method, DAG-Plan, generates a task graph, followed by task planning inference to iteratively generate nodes for each stage.
4.2 Experimental Results
Evaluation of Planning Effectiveness and Conciseness. As shown in Table 2, in the plan tests, DAG-Plan consistently outperformed both TP-S and TP-D, showcasing superior efficiency and robustness. DAG-Plan achieved a high success rate across all tasks, demonstrating its effectiveness in dual-arm manipulation. Notably, it maintained an impressive macro average success rate of 97.8% and exhibited a significant reduction in the required stages for task completion, averaging 5.67 stages. Furthermore, the completeness of the DAG generation is high, with only an incomplete DAG generated in task 8. By re-generating it through reflection, a complete DAG was obtained, increasing the success rate from 0.8 to 1.0. These results underscore the effectiveness of DAG-Plan in achieving task goals and its efficiency in execution. In contrast, TP-S, primarily focused on the single-arm plan, generally required more stages to complete tasks compared to TP-D and DAG-Plan. Although TP-S maintained a relatively high and consistent success rate, it was less efficient in stage minimization. Moreover, TP-D, relying on language models to generate dual-arm task plans, exhibited a significantly lower success rate (macro average SR of 44.5%), and often produced plans that were not executable in the physical environment. While theoretically capable of reducing the number of stages required for tasks, TP-D frequently encountered challenges related to coordination complexities and unrealistic task assignments. This highlights the superiority of DAG-Plan in effectively translating high-level plans into physical actions and navigating complexities in the environment with relative ease compared to TP-D. Additionally, we provide detailed plans of DAG-Plan and the baseline in Appendix C.
Table 2: Performance comparison on plan tests. We report the success rate and minimum stage of the 5 plans generated by the LLM for each task, and the macro average metrics for all tasks.
Scenario1 Scenario2
Task1 Task2 Task3 Task4 Task5 Task6
SR Stage SR Stage SR Stage SR Stage SR Stage SR Stage
TP-S $1.0$ $4$ $0.8$ $6$ $1.0$ $8$ $1.0$ $8$ $1.0$ $10$ $0.8$ $11$
TP-D $1.0$ $2$ $0.0$ $Fail$ $1.0$ $4$ $0.0$ $Fail$ $0.2$ $6$ $0.6$ $6$
DAG-Plan $1.0$ $2$ $1.0$ $4$ $1.0$ $4$ $0.8$ $6$ $1.0$ $6$ $1.0$ $6$
Scenario3 Macro Avg
Task7 Task8 Task9 SR Stage Stage Efficiency
SR Stage SR Stage SR Stage
TP-S $1.0$ $8$ $0.6$ $10$ $1.0$ $13$ $91.1\%$ $8.67\pm 2.54$ $100.0\%$
TP-D $0.6$ $5$ $0.6$ $7$ $0.0$ $Fail$ $44.5\%$ $6.33\pm 2.87$ $137.0\%$
DAG-Plan $1.0$ $5$ $1.0$ $7$ $1.0$ $11$ $97.8\%$ $5.67\pm 2.35$ $152.9\%$
Table 3: Performance comparison on physical simulation tests. We report the success rate and minimum time for each task with 10 trials, and the macro average metrics for all tasks.
Scenario1 Scenario2
Task1 Task2 Task3 Task4 Task5 Task6
SR Time SR Time SR Time SR Time SR Time SR Time
TP-S $0.7$ $37.1$ $0.7$ $59.0$ $0.7$ $79.6$ $1.0$ $84.1$ $0.3$ $105.5$ $0.3$ $114.4$
TP-D $1.0$ $18.6$ $0.0$ $Fail$ $0.6$ $40.1$ $0.0$ $Fail$ $0.0$ $Fail$ $0.0$ $Fail$
DAG-Plan $1.0$ $18.6$ $0.7$ $39.4$ $0.6$ $40.1$ $1.0$ $66.3$ $0.3$ $63.2$ $0.3$ $76.3$
Scenario3 Macro Avg
Task7 Task8 Task9 SR Time Execution Efficiency
SR Time SR Time SR Time
TP-S $0.5$ $74.0$ $0.1$ $105.2$ $0.2$ $136.2$ $49.9\%$ $88.3\pm 28.5$ $100.0\%$
TP-D $0.5$ $55.1$ $0.3$ $76.4$ $0.0$ $Fail$ $26.7\%$ $76.6\pm 35.5$ $115.3\%$
DAG-Plan $0.6$ $53.3$ $0.3$ $76.4$ $0.2$ $107.4$ $55.6\%$ $60.1\pm 24.5$ $147.0\%$
Evaluation with Physical Simulation. As shown in Table 3, the physical simulation tests provided further insights into the practical applicability and execution capabilities of our planning methods under more dynamic and realistic conditions. Here again, DAG-Plan demonstrated a balanced performance with a solid success rate and efficient execution times. We show the execution process of DAG-Plan in a physical simulation environment in Figure 4. Moreover, we provide detailed analysis and explanation of DAG-Plan and the baseline in Appendix D.
Compared to TP-D, DAG-Plan effectively translates high-level plans into feasible actions based on target object information and the robot’s current state under the guidance of a task graph. Both DAG-Plan and TP-S were able to complete $/9$ 9 tasks in the physical simulation tests. However, while TP-D passed the plan tests for $6/9$ tasks, it only completed $4/9$ in the physical simulation tests. In tasks 5 and 6, the plans generated by TP-D could not be executed due to real-world physical constraints. For example, in task 6, the plan of TP-D included “put cola bottle into refrigerator cooler" with left arm and “close microwave door" with right arm. However, since the refrigerator is on the right side of the microwave, the dual-arm robot could not cross its arms to execute this action. Ultimately, DAG-Plan achieved a 28.9% higher success rate in physical simulation tests compared to TP-D. This demonstrates that, both in high-level planning and low-level execution, the success rate of the dual-arm plans generated by TP-D is significantly inferior to those generated by DAG-Plan.
Regarding execution efficiency, the dual-arm plan allows for parallel execution of sub-tasks, resulting in higher efficiency. DAG-Plan’s execution efficiency was 47.0% higher than TP-S, as it maximized the parallelization of sub-tasks while ensuring the feasibility of the plan. In all tasks, DAG-Plan’s execution time was shorter than TP-S. Although TP-D had similar execution times to DAG-Plan for some tasks, its overall efficiency was low because $5/9$ tasks could not be executed. Consequently, TP-D’s efficiency was only 15.3% higher than TP-S.
Overall, the experimental results strongly support the adoption of DAG-Plan in complex robotic operations. DAG-Plan not only outperforms traditional single-arm and dual-arm planning approaches in terms of success rates and efficiency but also demonstrates significant robustness and reliability in translating plans into actionable steps in a physical context.
Figure 4: Head camera snapshots of the execution process of 2 example long-horizon tasks.
5 Conclusion
This work introduces DAG-Plan, which efficiently and accurately generates collaborative plans with LLMs for mobile dual-arm robots. DAG-Plan decomposes complex tasks into directed acyclic graph (DAG) with clear temporal relationships and iteratively selects feasible sub-tasks based on environmental observations during execution. The main contribution is the replacement of task sequences with a DAG and the dynamic adjustment of the planning according to the current situation, allowing dual-arm robots to flexibly utilize both arms for sub-task execution. We also conducted a dual-arm kitchen benchmark, providing a testing scenario for future long-horizon dual-arm works.
Limitations and Future Works. Currently, DAG-Plan relies on reinforcement learning for underlying skills, which may be inefficient for complex tasks. Our future efforts will aim to enhance skill learning modules, boosting automation and success rates for real-world applications.
Appendix A Implementation Details of Motion Planning and Reinforcement Learning
A.1 Embodied Platform
We used the Baxter robot from Rethink Robotics as the primary embodied platform shown in Figure 5. We configured its base in the simulation as a mobile base capable of translating on the plane but not rotating. The grippers of the Baxter robot were replaced with Robotiq 85 grippers to achieve more robust grasping.
Figure 5: Embodied platform of DAG-Plan.
A.2 Motion Planning
For motion planning, we use open-source implementations of RRT-connect [^2] https://github.com/haosulab/mplib. We manually set multiple candidate poses and use motion planning to find the pose that requires the shortest movement steps while avoiding collisions. For sub-tasks of the types occupy, tool use, and operate, we use motion planning to move the end-effector to within 0.15 cm of the target object, facing it as directly as possible. For release type sub-tasks, we adjust based on the opening direction of the target object. For high-precision sub-tasks that require both arms, such as “put pot into sink”, where motion planning alone struggles to maintain fixed distance and equal height for stable placement, we directly use reinforcement learning to learn the sub-task.
A.3 Reinforcement Learning
For reinforcement learning training, we use open-source implementations of PPO [^3] https://github.com/DLR-RM/stable-baselines3 algorithm, and list the hyper-parameters in Table 4. The observation space is the low-level state of the objects and robot and the images captured by 5 RGB-D cameras. The control mode of mobile base adopts velocity control. The control mode of the dual-arm adopts arm joint delta position control, which means the action space consists of the change of the arm joint position.
Table 4: Hyper-parameter of PPO algorithm.
Hyper-parameter Value
Discount factor $\gamma$ $0.99$
# of epochs per update $10$
[…]
# of environments $32$
Batch size $400$
Target KL divergence None
[…]
# of steps per update $4800$
Rollout steps per episode $100$
Corresponding rollout seconds per episode $4s$
Shape of observation space $(128,128,20)$Visual input
$95$State input
Shape of action space $18$
In this work, we use a set of reward primitives shown in Table 5. We use LLMs to combine reward primitives and specify the corresponding object and target object, with manual adjustments for corrections to generate dense rewards. The reward function and success condition for each sub-task are shown in Table 6. In the final reward, we normalize the reward functions of the subtasks corresponding to the left and right hands to 1 and assign additional rewards for successful states. The reward $r$ is represented as:
$r=r_{\text{right}}+r_{\text{left}}+r_{\text{side\_success}}\left\{\begin{aligned} &30,\text{if side success first time}\\ &0,\enspace\text{otherwise}\end{aligned}\right.+r_{\text{success}}\left\{\begin{aligned} &150-\text{steps},\text{if success}\\ &0,\quad\quad\quad\quad\enspace\text{otherwise}\end{aligned}\right.$
Table 5: Reward primitives.
Reward Primitives Description Pseudocode
$obj\_dis$ Move object close to target object $\frac{1}{2}(1-\frac{\text{dis}(obj,target)}{init\_dis}+1-\text{tanh}(\text{dis}(obj,target)))$
[…]
$obj\_grasped$ Grasp target object $1\enspace if\enspace\text{grasped}(target)\enspace else\enspace 0$
$obj\_in\_obj$ Put object into target object $1\enspace if\enspace\text{in}(obj,target)\enspace else\enspace 0$
$obj\_on\_obj$ Put object onto target object $1\enspace if\enspace\text{on}(obj,target)\enspace else\enspace 0$
$joint\_qpos$ Move joint to target joint position $\frac{\text{clip}(pos_{cur},pos_{min},pos_{max})-pos_{min}}{pos_{max}-pos_{min}}$
$ee\_pos$ Move end-effector to target object $\frac{1}{2}(1-\frac{\text{dis}(hand,target)}{init\_dis}+1-\text{tanh}(\text{dis}(hand,target)))$
$ee\_height$ Keep both end-effectors at same height $1-\frac{\text{min}((\text{abs}(z_{right}-z_{left})),0.05)}{0.05}$
$obj\_cut$ Cut the target object with object (knife) $\frac{1}{2}(obj\_grasped(obj)+obj\_on\_obj(obj,target))$
$obj\_pour$ Pour liquid in object into target object $\frac{1}{3}(obj\_grasped(obj)+obj\_horizon(obj)+obj\_above(obj,target))$
Table 6: Reward and success condition of Sub-tasks. Success condition is marked in bold.
Sub-task Pseudocode
$grasp<target\_obj>$ $1.0\times ee\_pos(target)+1.0\times\mathbf{obj\_grasped(target)}$
$pour<obj>into<target\_obj>$ $1.0\times\mathbf{obj\_pour(obj,target)}$
$cut<target\_obj>with<obj>$ $1.0\times\mathbf{obj\_cut(obj,target)}$
$put<obj>into<target\_obj>$ $3.0\times obj\_dis(obj,target)+1.0\times\mathbf{obj\_in(obj,target)}$
$put<pot>into<sink>$ $10.0\times ee\_both(sink)+3.0\times\mathbf{obj\_grasped(pot)}+1.0\times ee\_height+1.0\times\mathbf{obj\_in(pot,sink)}+1.0\times\mathbf{obj\_horizon(pot)}$
$put<obj>onto<target\_obj>$ $3.0\times obj\_dis(obj,target)+1.0\times\mathbf{obj\_on(obj,target)}$
$put<pot>onto<stove\_body>$ $10.0\times ee\_both(stove\_body)+3.0\times\mathbf{obj\_grasped(pot)}+1.0\times ee\_height+1.0\times\mathbf{obj\_on(pot,stove\_body)}+1.0\times\mathbf{obj\_horizon(pot)}$
$open<target\_obj>$ $3.0\times\mathbf{joint\_qpos(target\_joint,pos_{target})}+1.0\times obj\_dis(obj,target)+0.5\times obj\_grasped(obj,target)$
$close<target\_obj>$ $3.0\times\mathbf{joint\_qpos(target\_joint,pos_{target})}+1.0\times obj\_dis(obj,target)+0.5\times obj\_grasped(obj,target)$
$switch\enspace on<target\_obj>$ $3.0\times\mathbf{joint\_qpos(target\_joint,pos_{target})}+1.0\times obj\_dis(obj,target)+0.5\times obj\_grasped(obj,target)$
$switch\enspace off<target\_obj>$ $3.0\times\mathbf{joint\_qpos(target\_joint,pos_{target})}+1.0\times obj\_dis(obj,target)+0.5\times obj\_grasped(obj,target)$
Appendix B Dual-arm Kitchen Benchmark Details
B.1 Sub-tasks Action Capabilities
The action capabilities of sub-tasks in the Dual-arm Kitchen Benchmark are listed in Table 7.
Table 7: Sub-tasks action capabilities. Different actions correspond to the node types occupy, tool use, release, and operate.
Sub-task Description
$grasp<target\_obj>$ Grasp target object
$pour<obj>into<target\_obj>$ Pour the liquid in object into target object
$cut<target\_obj>with<obj>$ Cut target object which is food with object which is a knife
$put<obj>into<target\_obj>$ Put object in hand into target object
$put<obj>onto<target\_obj>$ Put object in hand onto target object
$open<target\_obj>$ Open target object, such as a door, a drawer floor, a faucet
$close<target\_obj>$ Close target object, such as a door, a drawer floor, a faucet
$switch\enspace on<target\_obj>$ Switch on target object which is an electric device
$switch\enspace off<target\_obj>$ Switch off target object which is an electric device
B.2 Full Sub-tasks List
The sub-tasks of each task in the Dual-arm Kitchen Benchmark are listed in Table 9.
Table 8: Sub-tasks of Dual-arm Kitchen Benchmark.
Index Instruction Sub-tasks
Task 1 Put the apple and bread onto the plate Grasp apple, Put apple onto plate, Grasp bread, Put bread onto plate
Task 2 Place the apple on the plate and toast the bread Grasp apple, Put apple onto plate, Grasp bread, Put bread into toaster body, Switch on toaster, Switch off toaster
Task 3 Juice the apple and toast the bread Grasp apple, Put apple into juicer container, Switch on juicer, Switch off juicer, Grasp bread, Put bread into toaster body, Switch on toaster, Switch off toaster
Task 4 Wash the cup and the bowl Open dishwasher door, Grasp cup, Put cup into dishwasher rack, Grasp bowl, Put bowl into dishwasher rack, Close dishwasher door, Switch on dishwasher, Switch off dishwasher
Task 5 Heat the soup and put the tin on the table Open refrigerator cooler door, Grasp bowl, Open microwave door, Put bowl into microwave body, Close microwave door, Switch on microwave, Switch off microwave, Grasp tin, Put tin onto table, Close refrigerator cooler door
Task 6 Heat the soup and pour a cup of cola Open refrigerator cooler door, Grasp bowl, Open microwave door, Put bowl into microwave body, Close microwave door, Switch on microwave, Switch off microwave, Grasp cola bottle, Pour cola bottle into cup, Put cola bottle into refrigerator cooler body, Close refrigerator cooler door
Task 7 Put the apple and the pear into the bowl Open cabinet left door, Grasp bowl, Put bowl onto table, Grasp apple, Put apple into bowl, Grasp pear, Put pear into bowl
Task 8 Cut the carrot and red bell pepper Open cabinet left door, Grasp knife, Grasp carrot, Put carrot onto cutting board, Cut carrot with knife, Grasp red bell pepper, Put red bell pepper onto cutting board, Cut red bell pepper, Put knife onto cutting board
Task 9 Make a pot of soup Open cabinet right door, Grasp pot, Put pot into sink, Open faucet, Close faucet, Grasp pot, Put pot onto stove body, Switch on stove, Grasp carrot, Put carrot into pot, Grasp red bell pepper, Put red bell pepper into pot, Switch off stove
B.3 Assets of Dual-arm Kitchen Benchmark
The assets of each task in the Dual-arm Kitchen Benchmark are listed in Table 9. These assets were sourced from Mobility-Partnet [^4] https://sapien.ucsd.edu/browse, BlenderKit [^5] https://www.blenderkit.com, and Sketchfab [^6] https://sketchfab.com. We have modified them and constructed the corresponding URDF models.
Table 9: Assets of Dual-arm Kitchen Benchmark.
Index Instruction Assets
Task 1 Put the apple and bread onto the plate table, apple, bread, juicer, toaster
Task 2 Place the apple on the plate and toast the bread table, apple, bread, juicer, toaster
Task 3 Juice the apple and toast the bread table, apple, bread, juicer, toaster
Task 4 Wash the cup and the bowl table, dishwasher, microwave, cup, bowl, refrigerator cooler
Task 5 Heat the soup and put the tin on the table table, dishwasher, microwave, tin, bowl, refrigerator cooler
Task 6 Heat the soup and pour a cup of cola table, dishwasher, microwave, cola bottle, bowl, refrigerator cooler
Task 7 Put the apple and the pear into the bowl table, sink, faucet, stove, cutting board, cabinet, shelf, knife, pot, apple, pear
Task 8 Cut the carrot and red bell pepper table, sink, faucet, stove, cutting board, cabinet, shelf, knife, pot, carrot, red bell pepper
Task 9 Make a pot of soup table, sink, faucet, stove, cutting board, cabinet, shelf, knife, pot, carrot, red bell pepper
Appendix C Planning Evaluation Analysis
To further illustrate the differences between our approach and the baselines, we analyzed the planning evaluation of TP-S, TP-D, and DAG-Plan for tasks 5 and 8. We validated the generated plans using Planning Domain Definition Language (PDDL) written by robotics experts for the corresponding tasks. The plans generated by TP-S and TP-D were executed sequentially in the PDDL environment. For DAG-Plan, due to the lack of environmental information in this evaluation, only the first check was conducted, and the first candidate node from the candidate list was selected for execution.
C.1 Planning Evaluation of Task 5
We provide the successful plan for each method in Figure 6 and the failed plan for TP-D in Figure 7. We further analyzed the failed plans of TP-D. TP-D made an error, incorrectly using the left hand, which was holding the tin, to close the refrigerator cooler door. In addition, TP-D is missing the sub-task of closing the microwave door.
C.2 Planning Evaluation of Task 8
We provide the successful plan for each method in Figure 8 and the failed plan for each method in Figure 9. We further analyzed the failed plans of each method. TP-S did not follow the rules in this task, failing to specify which hand to use for using the knife. TP-D made an error due to the complexity of dual-arm tasks, incorrectly using the left hand, which was not holding the knife, to cut. DAG-Plan did not follow the rules initially, generating an incomplete DAG, but after reflecting to LLMs for correction, it ultimately produced the same DAG as shown in Figure 8.
C.3 Failed Plan Analysis of TP-D
We further analyzed the failed plans of TP-D on the Dual-arm Kitchen Benchmark, as shown in Figure 10. It is evident that directly generating sequence dual-arm plans using LLMs often leads to errors, such as the robot holding an object in its hand while attempting to manipulate another object. This type of sequence temporal dependency works well for generating single-arm plans but struggles with the complexity of dual-arm planning, making it difficult to produce correct dual-arm plans. In contrast, DAG-Plan imposes strict constraints with occupy-release pair can be performed once the robot holds an object. This ensures that such errors are prevented, maintaining the conciseness of the plan.
Figure 6: Successful plan for each method of task 5.
Figure 7: Failed plan for each method of task 5. Errors are highlighted in red boxes.
Figure 8: Successful plan for each method of task 8.
Figure 9: Failed plan for each method of task 8. Errors are highlighted in red boxes.
Figure 10: Failed plan for TP-D of task 2, task 4, and task 9. Errors are highlighted in red boxes.
Appendix D Execution Process Analysis
D.1 Execution Process for DAG-Plan of Task 5
We provide the full execution process for DAG-Plan of task 5 in Figure 16.
(a) Stage 1: Open microwave door (left) and Open refrigerator cooler door (right)
(a) Stage 2: Grasp tin (left) and Grasp bowl (right)
(a) Stage 3: Put tin onto table (left) and Put bowl into microwave body (right)
(a) Stage 4: Close microwave door (left) and Close refrigerator cooler door (right)
(a) Stage 5: Switch on microwave (left)
(a) Stage 6: Switch off microwave (left)
Figure 16: Execution process for DAG-Plan of task 5.
D.2 Execution Process for DAG-Plan of Task 8
We provide the full execution process for DAG-Plan of task 5 in Figure 23.
(a) Stage 1: Grasp red bell pepper (left) and Grasp carrot (right)
(a) Stage 2: Put red bell pepper onto cutting board (left) and Put carrot onto cutting board (right)
(a) Stage 3: Open cabinet left door (right)
(a) Stage 4: Grasp knife (right)
(a) Stage 5: Cut carrot with knife (right)
(a) Stage 6: Cut red bell pepper with knife (right)
(a) Stage 7: Put knife onto cutting board (right)
Figure 23: Execution process for DAG-Plan of task 8.
D.3 Failed Execution Analysis of TP-D
Since TP-D generates dual-arm plans without the ability to adjust hand movements in real-time based on environmental information, some plans that pass PDDL validation still fail in the physical simulation environment. For instance, as shown in Figure 24, "Open refrigerator cooler door (left)" and "Open microwave door (right)" cannot be executed due to the hands crossing each other’s paths.
Figure 24: Failed Execution Process of TP-D.
Appendix E Full Prompts
E.1 Our Method: Prompt for DAG-Plan
{mdframed}
[backgroundcolor=azure,linewidth=0pt] You are a helpful assistant to plan a definite task into Directed Acyclic Graph (DAG) for a dual-arm robot.
I will give you the initial environment description:
task: …
obj_names: […]
obj_description: …
You must follow the following criteria:
1) Create a DAG based on the task name, considering the task complete when all nodes are executed. The task can be complex but should be broken into simple, reasonable nodes.
2) The node name must be selected in command API and follow the description.
3) When moving an object, confirm it clearly specify its destination, e.g., ‘put "obj_name" into/onto "obj_name"’
4) Synthesize the ’edge’ of a node, a list of node’s indices the precondition of this node. When generating edges, please match the current node with all other nodes to determine if there are any dependencies.
5) Your generated edges are not only the preconditions for the node to be executed, but also preconditions to complete the task objectives. For example, in task: fill cup with coffee, node: switch on coffee machine is a executed precondition for node: switch off coffee machine to execute, and additionally, node: put cup into coffee machine is also a task precondition for node: switch off coffee machine. If this condition is missing, switch off coffee machine may executed before put cup into coffee machine, the container will not filled with coffee, and the task cannot be completed.
6) Each node has a node type. "occupy" means the hand remains occupied after the node such as node: grasp "apple", and "release" means the object is released after this node such as node: put "apple" into "drawer", and "operate" means robot manipulate joint such as node: open "drawer", and "tool use" means robot utilize obj which handling in hand such as node: cut "apple" with "knife".
7) Each occupy node must have a corresponding release node which means the obj must holding in hand before put obj.
8) When two hands are needed to grasp or put an object, use a two-handed node directly instead of two one-handed nodes. For example, use node:grasp pot instead node: grasp pot_left_handle and node: grasp pot_right_handle.
9) Ensure the number of arms used matches the "arm_num" term. The "occupy" type nodes in the edge list should not exceed two.
10) Enclose all object references in quotes "". These must be in obj_names. Do not introduce new objects or remove any.
11) You need to generate as few nodes as possible in order to complete the task as quickly as possible while still being able to reach the end goal. Any node that is not relevant to the completion of the task should not be present.
12) After the graph is generated, create a task complete node.
List of commands of the API, all the object must selected in obj_names:
* The following command allows robot to open an <target-object> with a door, a drawer floor, a faucet:
  command syntax:
  -open <target-object>
  -close <target-object>
  node type: operate
* The following command allows robot to switch on or off an electric device. If the electric device has a door, such as a mircrowave with door, this command can only be selected when the door of electric device is closed:
  command syntax:
  - switch_on <target-object>
{mdframed}[backgroundcolor=azure, linewidth=0pt]   - switch_off <target-object>
  node type: operate
* The following command allows robot to put an <object> in or on a <target-object>, the <object> must holding in hand, the <target-object> should not holding in hand:
  command syntax:
  - put <object> into <target-object>
  - put <object> onto <target-object>
  node type: release
* The following command allows robot to pour a liquid from an <object> into an <target-object>. This command must be used to pour bottle of liquid:
  command syntax:
  - pour <object> into <target_object>
  node type: tool use
* The following command allows robot to cut a food which is <target-object> with an <object>. This <target-object> must on the cutting_board, after cutting the <target-object> keep same name before cutting, and still on cutting board not in hand:
  command syntax:
  -cut <target-object> with <object>
  node type: tool use
* The following command allows robot to take an <target-object>:
  command syntax:
  grasp <target-object>
  node type: occupy
Here’s an example input and response:
INPUT:
task: put apple into drawer
obj_names:  ["apple", "drawer_shell","drawer_floor", "table"]
obj_description:
    "apple":
        type: actor
        fixed: False
        spatial_relationship: "apple" is on "table".
        description: None
    "drawer":
        type: articulation
        fixed: True
        links: ["drawer_shell","drawer_floor","drawer_handle"]
        joints: ["drawer_joint"]
        spatial_relationship: "drawer" is on "table". "drawer" is closed now.
        description:
        "drawer_shell": This is the external part of the drawer, which cannot be manipulated but can be used to place items. The drawer shell is typically the external frame or casing of the drawer, providing structural support.
        "drawer_floor": This is the internal part of the drawer, containing a "drawer_handle" and "drawer_joint." The drawer floor can be opened or closed by manipulating the "drawer_handle."
         "drawer_handle": This is a handle located inside the drawer, and by operating it, the drawer can be opened or closed. Typically, the drawer handle is a component connected to the internal mechanism of the drawer.
        "drawer_joint": The parent link of this joint is "drawer_shell". The child link of this joint is "drawer_floor". It allows movement of the "drawer_floor" relative to the "drawer_shell," enabling the functionality of opening and closing the drawer.
{mdframed}[backgroundcolor=azure, linewidth=0pt]     "table":
        type: actor
        fixed: True
        spatial_relationship: "table" on the ground.
        description: None
RESPONSE:
nodes:
    node_1:
        name: open "drawer_floor"
        arm_num: 1
        edge: []
        type: operate
        description: Open the drawer to access inside.
    node_2:
        name: grasp "apple"
        arm_num: 1
        edge: []
        type: occupy
        description: Grasp the apple on the table.
    node_3:
        name: put "apple" into "drawer_floor"
        arm_num: 1
        edge: [1,2]
        type: release
        description: Put the apple into the drawer.
    node_4:
        name: task complete
        arm_num: 0
        edge: [3]
        type: complete
        description: Task complete after node_3 has been executed.
Now, let’s begin:
E.2 Baseline: Prompt for TP-S
{mdframed}
[backgroundcolor=azure, linewidth=0pt] You are a helpful assistant to plan a definite task into multi-satge for a dual-arm robot. To be on accuracy side, you can only use single arm or manipulate same object use dual arm each stage.
I will give you the initial environment description:
task: …
obj_names: […]
obj_description: …
You must follow the following criteria:
1) Break down the task into stages. The task can be complex but should be broken into simple, reasonable stages. The (side) should never be omitted.
2) The stage name must adhere the format: API (side). The API must select in command API and follow the description.
3) When moving an object, confirm it is grasped by hand and clearly specify its destination, e.g., ’put "obj_name" into/onto "obj_name"’.
4) Synethesis current hand empty or holding obj after stage has been executed. For example, right_hand: holding "apple", left_hand: empty.
{mdframed}[backgroundcolor=azure, linewidth=0pt] 5) The next stage you generate should accord to hand. For example, right_hand: holding "apple", left_hand: empty. The stage_next: put "apple" into "drawer", is reasonable then your right hand is empty, you can use your right_hand. The stage_next: grasp "pear" is wrong because your right hand state is holding "apple".
6) Each grasp stage should have a corresponding put stage which means a stage: grasp obj must have a corresponding stage: put obj.
7) When the obj in your hand has been used up, it should be returned to a fitting position to release your hands.
8) When two hands are needed to grasp or put an object, use a two-handed stage directly instead of two one-handed stages. For example, use stage: grasp pot (both) instead stage: grasp pot_left_handle (left) and stage_next: grasp pot_right_handle (right).
9) Enclose all noun references in quotes "". These must be in obj_names or ee_names. Do not introduce new objects or remove any.
10) You need to generate as few stages as possible in order to complete the task as quickly as possible while still being able to reach the end goal. Any stage that is not relevant to the completion of the task should not be present.
List of commands of the API, all the object must selected in obj_names:
* The following command allows robot to open an <target-object> with a door, a drawer floor, a faucet:
  command syntax:
  -open <target-object>
  -close <target-object>
* The following command allows robot to switch on or off an electric device. If the electric device has a door, such as a mircrowave with door, this command can only be selected when the door of electric device is closed:
  command syntax:
  - switch_on <target-object>
  - switch_off <target-object>
* The following command allows robot to put an <object> in or on a <target-object>, the <object> must holding in hand, the <target-object> should not holding in hand:
  command syntax:
  - put <object> into <target-object>
  - put <object> onto <target-object>
* The following command allows robot to pour a liquid from an <object> into an <target-object>. This command must be used to pour bottle of liquid:
  command syntax:
  - pour <object> into <target_object>
* The following command allows robot to cut a food which is <target-object> with an <object>. This <target-object> must on the cutting_board, after cutting the <target-object> keep same name before cutting, and still on cutting board not in hand:
  command syntax:
  -cut <target-object> with <object>
You should only respond in the format as described below and do not respond anything else:
RESPONSE FORMAT:
stages:
[…]
        name: The name of the stage.
        right_hand_state: Empty or holding sth.
        left_hand_state: Empty or holding sth.
{mdframed}[backgroundcolor=azure, linewidth=0pt] Here’s an example input and response:
INPUT:
task: put "apple" into "drawer_floor"
obj_names:  ["apple", "drawer_shell", "drawer_floor", "table"]
obj_description:
    "apple":
        type: actor
        fixed: False
        spatial_relationship: The "apple" is on the "table". The "apple" is close to "right_hand".
        description: None
    "drawer":
        type: articulation
        fixed: True
        links: ["drawer_shell","drawer_floor","drawer_handle"]
        joints:  ["drawer_joint"]
      spatial_relationship: This is a fixed articulation. The "drawer" is on the "table". The "drawer" is closed now. The "drawer" is close to "left_hand".
        description:
        "drawer_shell": This is the external part of the drawer, which cannot be manipulated but can be used to place items. The drawer shell is typically the external frame or casing of the drawer, providing structural support.
        "drawer_floor": This is the internal part of the drawer, containing a "drawer_handle" and "drawer_joint." The drawer floor can be opened or closed by manipulating the "drawer_handle."
        "drawer_handle": This is a handle located inside the drawer, and by operating it, the drawer can be opened or closed. Typically, the drawer handle is a component connected to the internal mechanism of the drawer.
        "drawer_joint": The parent link of this joint is "drawer_shell". The child link of this joint is "drawer_floor". It allows movement of the "drawer_floor" relative to the "drawer_shell," enabling the functionality of opening and closing the drawer.
    "table":
        type: actor
        fixed: True
        spatial_relationship: This is a fixed actor. The "drawer" and "apple" are on the "table".
        description: None
RESPONSE:
stages:
    stage_1:
        name: open "drawer_floor" (left)
        right_hand: empty
        left_hand: empty
        description: Open the drawer to access inside. This stage use left arm.
    stage_2:
        name: grasp "apple" (right)
        right_hand: holding "apple"
        left_hand: Grasp the apple on the table. This stage use right arm.
    stage_3:
        name: put "apple" into "drawer_floor" (right)
        right_hand: empty
        left_hand: empty
        description: Put the apple into the drawer use right arm. This stage use right arm.
Now, let’s begin:
E.3 Baseline: Prompt for TP-D
{mdframed}
[backgroundcolor=azure, linewidth=0pt]
You are a helpful assistant to plan a definite task into multi-satge for a dual-arm robot. To be on efficiency side, you can use dual arm each stage.
I will give you the initial environment description:
task: …
obj_names: […]
obj_description: …
You must follow the following criteria:
1) Break down the task into stages. The task can be complex but should be broken into simple, reasonable stages.
2) The stage name should adhere the format: "API (right) and API (left)". The API must select in command API and follow the description. Note that right API and left API is executing at the same time, without order.
3) If you need to use both hand too manipulate same object, the stage name should adhere the format: "API (both)".
4) When moving an object, confirm it is grasped by hand and clearly specify its destination, e.g., ’put "obj_name" into/onto "obj_name"’.
5) Synethesis current hand empty or holding obj after stage has been executed. For example, right_hand: holding "apple", left_hand: empty.
6) The next stage you generate should accroding to hand. For example, right_hand: holding "apple", left_hand: empty. The stage_next: put "apple" into "drawer", is reasonable then your right hand is empty, you can use your right_hand. The stage_next: grasp "pear" is wrong because your right hand state is holding "apple".
7) Each grasp stage should have a corresponding put stage which means a stage: grasp obj must have a corresponding stage: put obj.
8) When the obj in your hand has been used up, it should be returned to a fitting position to release your hands.
9) When two hands are needed to grasp or put an object, use a two-handed stage directly instead of two one-handed stages. For example, use stage: grasp pot (both) instead stage: grasp pot_left_handle (left) and grasp pot_right_handle (right).
10) Enclose all noun references in quotes "". These must be in obj_names or ee_names. Do not introduce new objects or remove any.
11) You need to generate as few stages as possible in order to complete the task as quickly as possible while still being able to reach the end goal. Any stage that is not relevant to the completion of the task should not be present.
12) Please try to take advantage of the dual-arm robot’s two-armed feature.
List of commands of the API, all the object must selected in obj_names:
* The following command allows robot to open an <target-object> with a door, a drawer floor, a faucet:
  command syntax:
  -open <target-object>
  -close <target-object>
* The following command allows robot to switch on or off an electric device. If the electric device has a door, such as a mircrowave with door, this command can only be selected when the door of electric device is closed:
  command syntax:
  - switch_on <target-object>
  - switch_off <target-object>
* The following command allows robot to put an <object> in or on a <target-object>, the <object> must holding in hand, the <target-object> should not holding in hand:
  command syntax:
  - put <object> into <target-object>
  - put <object> onto <target-object>
{mdframed}[backgroundcolor=azure, linewidth=0pt]
* The following command allows robot to pour a liquid from an <object> into an <target-object>. This command must be used to pour bottle of liquid:
  command syntax:
  - pour
<object> into <target_object>
* The following command allows robot to cut a food which is <target-object> with an <object>. This <target-object> must on the cutting_board, after cutting the <target-object> keep same name before cutting, and still on cutting board not in hand:
  command syntax:
  -cut <target-object> with <object>
* The following command allows robot to do nothing:
  command syntax:
  - None
You should only respond in the format as described below and do not respond anything else:
RESPONSE FORMAT:
stages:
[…]
        name: The name of the stage.
        right_hand_state: Empty or holding sth.
        left_hand_state: Empty or holding sth.
Here’s an example input and response:
INPUT:
task: put "apple" into "drawer_floor"
obj_names:  ["apple", "drawer_shell", "drawer_floor", "table"]
obj_description:
    "apple":
        type: actor
        fixed: False
        spatial_relationship: The "apple" is on the "table". The "apple" is close to "right_hand".
        description: None
    "drawer":
        type: articulation
        fixed: True
        links: ["drawer_shell","drawer_floor","drawer_handle"]
        joints:  ["drawer_joint"]
     spatial_relationship: This is a fixed articulation. The "drawer" is on the "table". The "drawer" is closed now. The "drawer" is close to "left_hand".
        description:
        "drawer_shell": This is the external part of the drawer, which cannot be manipulated but can be used to place items. The drawer shell is typically the external frame or casing of the drawer, providing structural support.
        "drawer_floor": This is the internal part of the drawer, containing a "drawer_handle" and "drawer_joint." The drawer floor can be opened or closed by manipulating the "drawer_handle."
         "drawer_handle": This is a handle located inside the drawer, and by operating it, the drawer can be opened or closed. Typically, the drawer handle is a component connected to the internal mechanism of the drawer.
         "drawer_joint": The parent link of this joint is "drawer_shell". The child link of this joint is "drawer_floor". It allows movement of the "drawer_floor" relative to the "drawer_shell," enabling the functionality of opening and closing the drawer.
    "table":
        type: actor
        fixed: True
{mdframed}[backgroundcolor=azure, linewidth=0pt]         spatial_relationship: This is a fixed actor. The "drawer" and "apple" are on the "table".
        description: None
RESPONSE:
stages:
    stage_1:
        name: grasp "apple" (right) and open "drawer_floor" (left)
        right_hand: holding "apple"
        left_hand: empty
       description: Grasp the apple on the table use right arm and open the drawer to access inside use left arm.
    stage_2:
        name: put "apple" into "drawer_floor" (right) and None (left)
        right_hand: empty
        left_hand: empty
        description: Put the apple into the drawer use right arm.
E.4 Insturction and Environment Description of Task5
{mdframed}
[backgroundcolor=azure, linewidth=0pt] task: heat the soup and put the tin on the table
constraints:
    1) do not need to take bowl out of microwave after heating
    2) close the door of refrigerator cooler after using
obj_names:["table", "dishwasher","dishwasher_rack", "dishwasher_door", "microwave","microwave_body", "microwave_door", "tin", "bowl", "refrigerator_cooler_body", "refrigerator_cooler_door"]
obj_description:
    "table":
        type: actor
        fixed: True
        spatial_relationship: The "table" is on the ground.
        description: None
    "dishwasher":
        type: articulation
        fixed: True
        links: ["dishwasher_shell", "dishwasher_rack", "dishwasher_door", "dishwasher_handle", "dishwasher_switch"
        joints: ["door_joint", "switch_joint"]
       spatial_relationship: The "dishwasher" is under the "table". The "dishwasher" is at the left part of the "table".
        description:
            "dishwasher_shell": The shell of the "dishwasher".
            "dishwasher_rack": The rack of the "dishwasher".
            "dishwasher_door": The door of the "dishwasher".
            "microwave_handle": The handle of the "dishwasher".
            "dishwasher_switch": The switch of the "dishwasher".
            "door_joint": The "door_joint" enables opening and closing of "dishwasher_door" by manipulating "microwave_handle".
            "switch_joint": The "switch_joint" enables and disables of "dishwasher" by switching off and switching on "dishwasher_switch". The "dishwasher_switch" can only be manipulated when "dishwasher_door" is closing.
    "microwave":
        type: articulation
        fixed: True
{mdframed}[backgroundcolor=azure, linewidth=0pt]         links: ["microwave_body", "microwave_door","microwave_handle","microwave_switch"]
       joints: ["door_joint","switch_joint"]
      spatial_relationship: The "microwave" is on the "table". The "microwave" is at the right part of the "table". The "microwave" at the left side of the "refrigerator".
        description:
            "microwave_body": The body of the "microwave".
            "microwave_door": The door of the "microwave".
            "microwave_handle": The handle of the "microwave".
            "microwave_switch": The switch of the "microwave".
            "door_joint": The "door_joint" enables opening and closing of "microwave_door" by manipulating "microwave_handle".
            "switch_joint": The "switch_joint" enables and disables of "microwave" by switching off and switching on "microwave_switch". The "microwave_switch" can only be manipulated when "microwave_door" is closing.
    "bowl":
        type: actor
        fixed: False
        spatial_relationship: The "bowl" is in the "refrigerator_cooler". The "bowl" is on the right side of the "tin".
        description: The "bowl" is filled with soup.
    "tin":
        type: actor
        fixed: False
        spatial_relationship: The "tin" is in the "refrigerator_cooler". The "tin" is on the left side of the "bowl".
        description: None
    "refrigerator_cooler":
        type: articulation
        fixed: True
   links: ["refrigerator_cooler_body", "refrigerator_cooler_door", "refrigerator_cooler_handle"]
        joints: ["door_joint"]
        spatial_relationship: The "refrigerator_cooler" is on the ground which at the right side the "table". The "refrigerator_cooler" at the right side of the "microwave". The "tin" and "bowl" both in the "refrigerator_cooler".
        description:
        "refrigerator_cooler_body": The body of the cooler floor of "refrigerator_cooler".         "refrigerator_cooler_door": The door of the cooler floor of "refrigerator_cooler".         "refrigerator_cooler_handle": The handle of the cooler floor door of "refrigerator_cooler".
        "door_joint": The "door_joint" enables opening and closing of "refrigerator_cooler_door" by manipulating "refrigerator_cooler_handle".
E.5 Insturction and Environment Description of Task8
{mdframed}
[backgroundcolor=azure, linewidth=0pt] task: cut carrot and red bell pepper
constraints:
    1) After cutting carrot and red bell pepper, put the knife onto the cutting board
obj_names:["table", "sink", "faucet", "stove", "stove_body", "cutting_board", "cabinet_body", "cabinet_left_door", "cabinet_right_door", "knife_rack", "shelf_bottom_floor", "shelf_top_floor", "knife", "bowl", "pot", "carrot", "red_bell_pepper"]
obj_description:
    "table":
        type: actor
        fixed: True
        spatial_relationship: The "table" is on the ground.
        description: None
    "sink":
{mdframed}[backgroundcolor=azure, linewidth=0pt]         type: actor
        fixed: True
        spatial_relationship: The "sink" is on right side of the table.
        description: A sink on the right side of the table. It offers a deep basin, can put the pot into it for filling water it by faucet.
    "faucet":
        type: articulation
        fixed: True
        links: ["faucet_body", "faucet_handle"]
        joints: ["faucet_joint"]
        spatial_relationship: The "faucet" is on the "table". The "faucet" is at the back of the sink.
        description:
        "faucet_body": "When pot in the sink and faucet is openning, pot can be filled with water."
        "faucet_handle": "The faucet handle of faucet."
        "faucet_joint": "The "faucet_joint" enables opening and closing of "faucet" by oepning and closing "faucet_handle". It provides flexibility in directing water flow."
    "stove":
        type: articulation
        fixed: True
        links: ["stove_body", "fire_knob"]
        joints: ["stove_joint"]
        spatial_relationship: The "stove" is on the "table". The "stove" is at the left side of "sink".
        description:
        "stove_body": "It allows which put on it for versatile cooking options."
        "fire_knob": "A knob in front of the stove. The fire knob provides easy ignition control."
        "stove_joint": "The "stove_joint" enables switching on and switching off of "stove" by manipulate "fire_knob"."
    "cutting_board":
        type: actor
        fixed: True
        spatial_relationship: The "cutting_board" is on the table.
        description: "A cutting board on the left front side of the table. You can’t directly cut food in your hand. Any food must be put on the cutting board first before cutting."
    "cabinet":
        type: articulation
        fixed: True
        links: ["cabinet_body", "cabinet_door_left", "cabinet_left_handle", "cabinet_door_right", "cabinet_right_handle", "knife_rack"]
        joints: ["cabinet_right_joint", "cabinet_left_joint"]
      spatial_relationship: The "tool_cabinet" is above "table". The "cabinet_door_left" and "cabinet_door_right" is close.
        description:
         "cabinet_body": "It provides three storage space for bowl, knives and pot from left to right. The "bowl", "knife", and "knife_rack" is in the left part of "cabinet". The "pot" is in the right part of "tool_cabinet"."
         "cabinet_left_door": "The left door of the cabinet, only it opened can access to "knife_rack"."
        "cabinet_left_handle": "The left door is equipped handle for convenient opening and closing."
        "cabinet_right_door": "The right door of the cabinet, only it opened can access to the right half of the cabinet."
        "cabinet_right_handle": "The right door is equipped handle for convenient opening and closing."
         "knife_rack": "The knife rack inside the left part of cabinet. It can securely hold kitchen knives. It has three slots for knives."
{mdframed}[backgroundcolor=azure, linewidth=0pt]         "cabinet_right_joint": "The "cabinet_right_joint" enables opening and closing of "cabinet_door_right" by manipulating "cabinet_right_handle"."
        "cabinet_left_joint": "The "cabinet_left_joint" enables opening and closing of "cabinet_door_left" by manipulating "cabinet_left_handle"."
    "shelf":
        type: articulation
        fixed: True
        links: ["shelf_bottom_floor", "shelf_top_floor"]
        joints: None
      spatial_relationship: The "shelf" is on the table. The "shelf" at the back side of "cutting_board".
        description:
            "shelf_bottom_floor": "The shelf has two floor. The "shelf_bottom" is the bottom floor of "shelf"."
            "shelf_top_floor": "The "shelf_top" is the top floor of "shelf"."
    "knife":
        type: actor
        fixed: False
        spatial_relationship: "knife" is in "knife_rack".
        description: None
    "bowl":
        type: actor
        fixed: False
        spatial_relationship: "bowl" is in "cabinet_body". "bowl" is at the left of "knife_rack".
        description: None
    "pot":
        type: articulation
        fixed: False
        links: ["pot_body", "pot_left_handle", "pot_right_handle"]
        joints: None
        spatial_relationship: "pot" is in "cabinet_body".
        description:
            "pot_body": "The body of the "pot". It has a lot of mass and if you want to move it, you need to hold the handles with both hand."
            "pot_left_handle": The left handle of the "pot".
            "pot_right_handle": The right handle of the "pot".
    "carrot":
        type: actor
        fixed: False
        spatial_relationship: "carrot" is on "shelf_bottom_floor".
        description: None
    "red_bell_pepper":
        type: actor
        fixed: False
        spatial_relationship: "red_bell_pepper" is on "shelf_bottom_floor".
        description: None
Appendix F Example Outputs
F.1 Our method: DAG-Plan Output of Task 5
{mdframed}
[backgroundcolor=azure, linewidth=0pt] nodes:
node_1:
name: open "refrigerator_cooler_door"
arm_num: 1
edge: []
type: operate
description: Open the refrigerator cooler door to access the contents inside.
node_2:
name: grasp "bowl"
arm_num: 1
edge: [1]
type: occupy
description: Grasp the bowl containing soup from the refrigerator cooler.
node_3: name: open "microwave_door"
arm_num: 1
edge: []
type: operate
description: Open the microwave door to place the bowl inside.
node_4: name: put "bowl" into "microwave"
arm_num: 1
edge: [2, 3]
type: release
description: Place the bowl inside the microwave for heating.
node_5: name: close "microwave_door"
arm_num: 1
edge: [4]
type: operate
description: Close the microwave door to start the heating process.
node_6:
name: switch_on "microwave"
arm_num: 1
edge: [5]
type: operate
description: Switch on the microwave to heat the soup in the bowl.
node_7:
name: switch_off "microwave"
arm_num: 1
edge: [6]
type: operate
description: Switch off the microwave after the soup is heated.
node_8:
name: grasp "tin"
arm_num: 1
edge: [1]
type: occupy
description: Grasp the tin from the refrigerator cooler.
node_9:
name: put "tin" onto "table"
arm_num: 1
edge: [8]
type: release
description: Place the tin onto the table.
{mdframed}[backgroundcolor=azure, linewidth=0pt] node_10:
name: close "refrigerator_cooler_door"
arm_num: 1
edge: [2, 8]
type: operate
description: Close the refrigerator cooler door after taking out the bowl and tin.
node_11:
name: task complete
arm_num: 0
edge: [7, 9, 10]
type: complete
description: Task complete after heating the soup and placing the tin on the table, and closing the refrigerator cooler door.
F.2 Baseline: TP-S Output of Task 5
{mdframed}
[backgroundcolor=azure, linewidth=0pt] stages:
stage_1:
name: open "refrigerator_cooler_door" (left)
right_hand_state: empty
left_hand_state: empty
description: Open the refrigerator cooler door to access the items inside. This stage uses the left arm.
stage_2:
name: grasp "tin" (right)
right_hand_state: holding "tin"
left_hand_state: empty
description: Grasp the tin from the refrigerator cooler. This stage uses the right arm.
stage_3: name: put "tin" onto "table" (right)
right_hand_state: empty
left_hand_state: empty
description: Place the tin onto the table. This stage uses the right arm.
stage_4:
name: grasp "bowl" (right)
right_hand_state: holding "bowl"
left_hand_state: empty
description: Grasp the bowl filled with soup from the refrigerator cooler. This stage uses the right arm.
stage_5:
name: open "microwave_door" (left)
right_hand_state: holding "bowl"
left_hand_state: empty
description: Open the microwave door to place the bowl inside. This stage uses the left arm.
stage_6:
name: put "bowl" into "microwave_body" (right)
right_hand_state: empty
left_hand_state: empty
description: Place the bowl inside the microwave for heating. This stage uses the right arm.
stage_7:
name: close "microwave_door" (left)
right_hand_state: empty
left_hand_state: empty
description: Close the microwave door after placing the bowl inside. This stage uses the left arm.
{mdframed}[backgroundcolor=azure, linewidth=0pt] stage_8:
name: switch_on "microwave" (right)
right_hand_state: empty
left_hand_state: empty
description: Switch on the microwave to start heating the soup. This stage uses the right arm.
stage_9:
name: switch_off "microwave" (right)
right_hand_state: empty
left_hand_state: empty
description: Switch off the microwave after the soup is heated. This ensures safety by keeping the appliance off after use. This stage uses the right arm.
stage_10:
name: close "refrigerator_cooler_door" (left)
right_hand_state: empty
left_hand_state: empty
description: Close the refrigerator cooler door after retrieving and returning items. This stage uses the left arm.
F.3 Baseline: TP-D Output of Task 5
{mdframed}
[backgroundcolor=azure, linewidth=0pt]
stages:
stage_1:
name: open "refrigerator_cooler_door" (left) and open "microwave_door" (right)
right_hand_state: empty
left_hand_state: empty
description: Open the refrigerator cooler door with left hand and microwave door with right hand to access the bowl and microwave.
stage_2:
name: grasp "bowl" (left) and grasp "tin" (right)
right_hand_state: holding "tin"
left_hand_state: holding "bowl"
description: Grasp the bowl with soup from the refrigerator with left hand and the tin with right hand.
stage_3:
name: put "bowl" into "microwave_body" (left) and put "tin" onto "table" (right)
right_hand_state: empty
left_hand_state: holding "bowl"
description: Place the bowl inside the microwave with left hand and put the tin on the table with right hand.
stage_4:
name: close "refrigerator_cooler_door" (right) and switch_on "microwave" (left)
right_hand_state: empty
left_hand_state: empty
description: Close the refrigerator cooler door using the right hand and switch on the microwave with the left hand to start heating the soup.
stage_5:
name: switch_off "microwave" (left) and None (right)
right_hand_state: empty
left_hand_state: empty
description: Turn off the microwave after heating the soup to ensure safety, with the left hand.
F.4 Our method: DAG-Plan Output of Task 8
{mdframed}
[backgroundcolor=azure, linewidth=0pt] nodes:
node_1:
name: open "cabinet_left_door"
arm_num: 1
edge: []
type: operate
description: Open the left door of the cabinet to access the knife rack.
node_2:
name: grasp "knife"
arm_num: 1
edge: [1]
type: occupy
description: Grasp the knife from the knife rack.
node_3:
name: grasp "carrot"
arm_num: 1
edge: []
type: occupy
description: Grasp the carrot from the shelf bottom floor.
node_4:
name: put "carrot" onto "cutting_board" arm_num: 1
edge: [3]
type: release
description: Place the carrot onto the cutting board to prepare for cutting.
node_5:
name: cut "carrot" with "knife"
arm_num: 1
edge: [2, 4]
type: tool use
description: Cut the carrot using the knife while it is on the cutting board.
node_6:
name: grasp "red_bell_pepper"
arm_num: 1
edge: []
type: occupy
description: Grasp the red bell pepper from the shelf bottom floor.
node_7:
name: put "red_bell_pepper" onto "cutting_board"
arm_num: 1
edge: [6]
type: release
description: Place the red bell pepper onto the cutting board to prepare for cutting.
node_8:
name: cut "red_bell_pepper" with "knife"
arm_num: 1
edge: [2, 7]
type: tool use
description: Cut the red bell pepper using the knife while it is on the cutting board.
node_9:
name: put "knife" onto "cutting_board"
arm_num: 1
edge: [5, 8]
type: release
description: Place the knife back onto the cutting board after use.
{mdframed}[backgroundcolor=azure, linewidth=0pt] node_10:
name: task complete
arm_num: 0
edge: [9]
type: complete
description: Task complete after all cutting and placements are done.
F.5 Baseline: TP-S Output of Task 8
{mdframed}
[backgroundcolor=azure, linewidth=0pt] stages:
stage_1:
name: open "cabinet_left_door" (left)
right_hand_state: empty
left_hand_state: empty
description: Open the left door of the cabinet to access the knife rack.
stage_2:
name: grasp "knife" (right)
right_hand_state: holding "knife"
left_hand_name: empty
description: Grasp the knife from the knife rack.
stage_3:
name: close "cabinet_left_door" (left)
right_hand_state: holding "knife"
left_hand_state: empty
description: Close the left door of the cabinet after taking out the knife.
stage_4:
name: grasp "carrot" (left)
right_hand_state: holding "knife"
left_hand_state: holding "carrot"
description: Grasp the carrot from the shelf bottom floor.
stage_5:
name: put "carrot" onto "cutting_board" (left)
right_hand_state: holding "knife"
left_hand_state: empty
description: Place the carrot onto the cutting board to prepare for cutting.
stage_6:
name: cut "carrot" with "knife"
right_hand_state: holding "knife"
left_hand_state: empty
description: Cut the carrot on the cutting board using the knife.
stage_7:
name: grasp "red_bell_pepper" (left)
right_hand_state: holding "knife"
left_hand_state: holding "red_bell_pepper"
description: Grasp the red bell pepper from the shelf bottom floor.
stage_8:
name: put "red_bell_pepper" onto "cutting_board" (left)
right_hand_state: holding "knife"
left_hand_state: empty
description: Place the red bell pepper onto the cutting board to prepare for cutting.
stage_9:
name: cut "red_bell_pepper" with "knife"
right_hand_state: holding "knife"
left_hand_state: empty
description: Cut the red bell pepper on the cutting board using the knife.
{mdframed}[backgroundcolor=azure, linewidth=0pt] stage_10:
name: put "knife" onto "cutting_board" (right)
right_hand_state: empty
left_hand_state: empty
description: Place the knife back onto the cutting board after cutting the vegetables.
F.6 Baseline: TP-D Output of Task 8
{mdframed}
[backgroundcolor=azure, linewidth=0pt]
stages:
stage_1:
name: grasp "carrot" (right) and grasp "red_bell_pepper" (left)
right_hand_state: holding "carrot"
left_hand_state: holding "red_bell_pepper"
description: The robot grasps the "carrot" with its right hand and the "red_bell_pepper" with its left hand from the "shelf_bottom_floor".
stage_2:
name: put "carrot" onto "cutting_board" (right) and put "red_bell_pepper" onto "cutting_board" (left)
right_hand_state: empty
left_hand_state: empty
description: The robot places both the "carrot" and "red_bell_pepper" onto the "cutting_board" to prepare them for cutting.
stage_3:
name: open "cabinet_left_door" (left) and None (right)
right_hand_state: empty
left_hand_state: empty
description: The robot opens the left door of the "cabinet" to access the "knife_rack".
stage_4:
name: grasp "knife" (right) and None (left)
right_hand_state: holding "knife"
left_hand_state: empty
description: The robot grasps the "knife" from the "knife_rack".
stage_5:
name: cut "carrot" with "knife" (right) and cut "red_bell_pepper" with "knife" (left)
right_hand_state: holding "knife"
left_hand_state: empty
description: The robot uses the "knife" to cut both the "carrot" and the "red_bell_pepper" on the "cutting_board".
stage_6:
name: put "knife" onto "cutting_board" (right) and None (left)
right_hand_state: empty
left_hand_state: empty
description: After cutting, the robot places the "knife" back onto the "cutting_board" as per the task constraints.

### v3

@@
 Dual-arm robots offer enhanced versatility and efficiency over single-arm counterparts by enabling concurrent manipulation of multiple objects or cooperative execution of tasks using both arms.
-However, effectively coordinating the two arms for complex long-horizon tasks remains a significant challenge.
-Existing task planning methods predominantly focus on single-arm robots or rely on predefined bimanual operations, failing to fully leverage the capabilities of dual-arm systems.
+However, the coordination of dual-arm systems for long-horizon tasks continues to pose significant challenges, stemming from the intricate temporal and spatial dependencies among sub-tasks, necessitating intelligent decisions regarding the allocation of actions between arms and their optimal execution order.
+Existing task planning methods predominantly focus on single-arm robots or rely on predefined bimanual operations to use large language models (LLMs) generate task sequence with linear temporal dependency, failing to fully leverage the capabilities of dual-arm systems.
 To address this limitation, we introduce DAG-Plan, a structured task planning framework tailored for dual-arm robots.
-DAG-Plan harnesses large language models (LLMs) to decompose intricate tasks into actionable sub-tasks represented as nodes within a directed acyclic graph (DAG).
+DAG-Plan harnesses LLMs to decompose intricate tasks into actionable sub-tasks represented as nodes within a directed acyclic graph (DAG).
 Critically, DAG-Plan dynamically assigns these sub-tasks to the appropriate arm based on real-time environmental observations, enabling parallel and adaptive execution.
-We evaluate DAG-Plan on the novel Dual-Arm Kitchen Benchmark, comprising 9 sequential tasks with 78 sub-tasks and 26 objects.
-Extensive experiments demonstrate the superiority of DAG-Plan over directly using LLM to generate plans, achieving nearly 50% higher efficiency compared to the single-arm task planning baseline and nearly double the success rate of the dual-arm task planning baseline.
-Keywords: Dual-arm Robots, Task Planning, Large Language Models
-1 Introduction
-Achieving effective bimanual coordination in robotics is challenging due to the complexities of dual-arm operations, requiring precise spatial and temporal coordination [1, 2].
-While humans effortlessly coordinate their hands in daily tasks, replicating such coordination in robots presents significant challenges for both planning and learning-based methods.
-Traditionally, planning-based methods have focused on motion planning, employing hand-designed primitives to manage the movements of two robotic arms [3, 4].
-However, these methods often fall short in dynamic or intricate environments as they lack the flexibility required for adaptive task execution.
-In contrast, learning-based strategies, such as Reinforcement Learning (RL) and Imitation Learning (IL), provide more adaptive solutions by enabling robots to learn control policies from either human-designed rewards [5, 6] or human demonstrations [7, 8].
-However, these methods struggle to generalize to zero-shot scenarios, where robots need to execute tasks without prior specific training.
-Large language models (LLMs) have emerged as powerful tools endowed with extensive knowledge and sophisticated reasoning abilities [9, 10, 11, 12].
-By systematically breaking down tasks into actionable sub-tasks and leveraging their commonsense knowledge and implicit reasoning capabilities, LLMs empower robots to effectively adapt to new scenarios and tasks [13, 14, 15, 16, 17, 18].
-However, existing planning methods are primarily applied to single-arm robots, focusing on using one arm to perform skills.
+We evaluate DAG-Plan on the Dual-Arm Kitchen Benchmark, comprising 5 sequential tasks with 44 sub-tasks.
+Extensive experiments demonstrate the superiority of DAG-Plan over directly using LLM to generate linear task sequence, achieving 52.8% higher efficiency compared to the single-arm task planning and 48% higher success rate of the dual-arm task planning.
+Compared to iterative methods, DAG-Plan improving execution efficiency 84.1% due to its fewer query time.
+More demos and information are available on https://sites.google.com/view/dag-plan.
+Index Terms:
+Dual-arm Robots, Task Planning, LLMs
+I Introduction
+Achieving effective bimanual coordination in robotics is challenging due to the complexities and long-horizon of dual-arm operations, requiring precise spatial and temporal coordination [1, 2].
+While humans effortlessly coordinate their hands in daily long-horizon tasks, replicating such coordination in robots presents significant challenges.
+Traditional method employed hand-designed primitives to manage the movements of dual robotic arms [3, 4].
+Whereas these methods often fall short in long-horizon task as they lack the flexibility required for adaptive task execution.
+Large language models (LLMs) have emerged as powerful tools endowed with extensive knowledge and sophisticated reasoning abilities [5, 6].
+By systematically breaking down tasks into actionable sub-tasks and leveraging their commonsense knowledge and implicit reasoning capabilities, LLMs empower robots to effectively adapt to long-horizon scenarios and tasks in the wild [7, 8].
+Existing LLMs planning methods are primarily applied to single-arm robots, focusing on using one arm to perform skills.
 While some studies [9, 10] have employed dual-arm robots as test platforms, these still engage only one arm at a time or rely on predefined bimanual tasks, leading to inefficiencies.
-In this work, we propose the development of dual-arm cooperative planning using LLMs.
-This approach facilitates parallel task execution, thereby significantly improving efficiency.
-Directly generating dual-arm planning schemes as subgoal sequences with LLMs faces significant challenges, primarily two gaps: 1) Dual-arm collaboration allows multiple sub-tasks to be executed simultaneously, making the temporal dependencies between them highly complex.
-Previous approaches, which employed linear temporal dependency lists for task sequencing, have proven to be inefficient for planning and execution; 2) The non-interactivity with the environment, as the execution order of the task sequence and the side of the executing arm are fixed, making it impossible to choose executable and cost-effective sub-tasks based on the environment and robot state during execution.
+Several multi-agent planning methods  [11, 12, 13] have been proposed, which assign tasks to each robot through independent task decomposition or interactions among multiple robots, significantly enhancing performance in multi-robot collaborative tasks.
+These methods share a common characteristic: they all utilize LLMs to generate task sequences with linear temporal dependencies.
+Generating dual-arm planning schemes as task sequences with LLMs faces significant challenges, primarily three gaps: 1) Dual-arm collaboration allows multiple sub-tasks to be executed simultaneously, making the temporal dependencies between them highly complex.
+Previous approaches, which employed linear temporal dependency lists for task sequencing, have proven to be inefficient for planning and execution; 2) The non-interactivity with the environment, as the execution order of the task sequence and the side of the executing arm are fixed in non-iterative method, making it impossible to choose executable and cost-effective sub-tasks based on the environment and robot state during execution; 3) The iterative approach, while capable of interacting with the environment and achieving a higher success rate compared to non-iterative methods, necessitates querying the LLM for each decision, resulting in significant token consumption and time costs.
 To address these challenges, we introduce DAG-Plan, a structured task planning framework that leverages the capabilities of LLMs.
@@
 DAG-Plan checks the feasibility and calculates the cost of combinations of left and right arm candidate nodes based on the environmental state, adaptively executing sub-tasks that are easier and closer to perform.
-Our main contributions are summarized as follows: 1) We present DAG-Plan, an efficient cooperative task planning framework for mobile dual-arm robots.
-This framework represents decomposed sub-tasks as a DAG and dynamically assigns these sub-tasks to the appropriate arm based on the real-time environment state; 2) We conduct the Dual-arm Kitchen Benchmark based on Sapien engine [21], which consists of 9 long-horizon tasks with 78 sub-tasks and 26 assets and objects.
-This Benchmark is manually constructed by robotic experts and includes plan tests and physical simulation tests; 3) Extensive experiments on the Dual-arm Kitchen Benchmark show that DAG-Plan significantly outperforms other methods.
-DAG-Plan achieves nearly a 50% increase in efficiency over the baseline by directly employing LLMs to generate the single-arm plan.
-Compared to the baseline, which directly utilizes LLMs for dual-arm plans, DAG-Plan demonstrates nearly double the success rate.
+Our main contributions are summarized as follows:
+  • We identify the limitations of planning methods which generate linear task sequences when applied to dual-arm robots, particularly their inefficiency in handling complex temporal dependencies and their inability to adaptively interact with the environment during task execution.
+  • We present DAG-Plan, an efficient cooperative task planning framework for dual-arm robots.
+This framework represents decomposed sub-tasks as a DAG with LLMs and dynamically assigns sub-tasks to the appropriate arm based on the real-time environment state.
+  • Extensive experiments on the Dual-arm Kitchen Benchmark show that DAG-Plan significantly outperforms other methods.
+DAG-Plan achieves 52.8% increase in efficiency over the baseline by directly employing LLMs to generate the single-arm plan.
+Compared to the baseline, which directly utilizes LLMs for dual-arm plans, DAG-Plan demonstrates 48% higher success rate.
+Compared to iterative methods, DAG-Plan improving execution efficiency 84.1% due to its fewer query time.
+II Related Works
+II-A Task Planning with LLMs
+LLMs are increasingly being used to generate sequences of executable actions that enable an agent to achieve goals represented in natural language.
+Previous studies have successfully utilized the commonsense and in-context learning capabilities of pre-trained LLMs to create executable plans for embodied agents [14, 9, 10, 15, 11, 12, 13].
+However, most of these studies have been applied to single-arm robots, only requiring consideration of executing a single-arm action at one timestep.
+Additionally, these studies [9, 11] explicitly use the operational rules of the environment as input, whereas we only provide an environment description, relying on the LLM’s inherent world modeling capability.
+Multi-agents planning is the closest to dual-arm planning.
+Twostep [11] decomposes task into two independent tasks.
+RoCo [12] and DABICO [13] assign each robot with a LLM to discuss and collectively reason task strategies.
+These methods either fail to consider the constraints of dual-arm robot or are limited by the linear dependencies of task sequence.
+II-B Dual-arm Robot Manipulation
+Dual-arm coordination has made progress in industrial [1, 2] and agricultural scenarios [16, 17] with fixed process operations and domestic settings [18, 19] with single-skill operations.
+With the development of methods such as motion planning [17, 19], foundation model [20, 21], reinforcement learning [22, 23], and imitation learning [24, 25], dual-arm robots can perform many human-like operations at the skill level.
+However, they still lack the ability to autonomously execution in long-horizon task.
+The commonsense and contextual learning capabilities of pre-trained LLMs make it possible for dual-arm robots to autonomously execution.
+II-C Structured Task Decomposition (STD)
+STD involves breaking down a complex task into a DAG.
+Previous methods for STD, such as Crowd-Sourced STD [26, 27] and Query-based STD [28, 29], were limited by data availability.
+However, LLMs contain extensive real-world commonsense knowledge, offering new approaches for STD.
+TaskLAMA [30] has conducted detailed research on structured task decomposition using LLMs, demonstrating that LLMs can decompose real-world tasks into task graphs with temporal dependencies.
+In this work, we explore the use of DAG to address issues in dual-arm robot sequential planning with low execution efficiency.
+III Methods
+We utilize LLMs to generate a DAG, where each task for a dual-arm robot is represented as a node.
+The directed edges between these nodes are crucial as they establish a clear and mandatory sequence of tasks, dictating the order in which tasks must be performed.
+This ensures that dependencies are meticulously adhered to, allowing for efficient task execution.
+The robot dynamically assesses the state of its environment and the status of its arms to select next optimal tasks from the graph.
+This ongoing selection process prioritizes activities, keeping both arms continuously engaged.
+An overview of DAG-Plan pipeline is illustrated in Figure 1.
 Figure 1: An overview of DAG-Plan.
@@
 The framework then evaluates all candidate combinations for feasibility and cost.
-DAG-Plan selects the nodes with the lowest cost and employs motion planning and reinforcement learning for execution.
+DAG-Plan selects the nodes with the lowest cost and employs skill in library for execution.
 DAG-Plan updates the graph, iterating inference until the DAG is fully executed.
-2 Related Works
-Task Planning with LLMs.
-LLMs are increasingly being used to generate sequences of executable actions that enable an agent to achieve goals represented in natural language.
-Previous studies have successfully utilized the commonsense and in-context learning capabilities of pre-trained LLMs to create executable plans for embodied agents [22, 23, 19, 24, 25, 20, 26, 27, 28].
-However, most of these studies have been applied to single-arm robots, only requiring consideration of executing a single-arm action at one timestep.
-Although LLM+P [19] and CoPAL [20] both have designed the tasks for dual-arm robots.
-The plan generated by LLM+P and CoPAL still can not manipulate two different objects with each arm at the same time causing low execution efficiency.
-Additionally, these studies [19, 27, 28] explicitly use the operational rules of the environment as input, whereas we only provide an environment description, relying on the LLM’s inherent world modeling capability.
-Dual-arm Robot Manipulation.
-Dual-arm coordination has made progress in industrial [1, 2] and agricultural scenarios [29, 30] with fixed process operations and domestic settings [31, 32] with single-skill operations.
-With the development of methods such as motion planning [30, 32], reinforcement learning [33, 34], and imitation learning [35, 36], dual-arm robots can perform many human-like operations at the skill level.
-However, they still lack the ability to autonomously plan in zero-shot complex scenarios.
-The commonsense and contextual learning capabilities of pre-trained LLMs make it possible for dual-arm robots to autonomously plan in zero-shot complex scenarios.
-Structured Task Decomposition (STD).
-STD involves breaking down a complex task into a DAG.
-Previous methods for STD, such as Crowd-Sourced STD [37, 38] and Query-based STD [39, 40], were limited by data availability.
-However, LLMs contain extensive real-world commonsense knowledge, offering new approaches for STD.
-TaskLAMA [41] has conducted detailed research on structured task decomposition using LLMs, demonstrating that LLMs can decompose real-world tasks into task graphs with temporal dependencies.
-In this work, we explore the use of DAG to address issues in dual-arm robot sequential planning with low execution efficiency.
-3 Method
-We utilize LLMs to generate a DAG, where each task for a dual-arm robot is represented as a node.
-The directed edges between these nodes are crucial as they establish a clear and mandatory sequence of tasks, dictating the order in which tasks must be performed.
-This ensures that dependencies are meticulously adhered to, allowing for efficient task execution.
-The robot dynamically assesses the state of its environment and the status of its arms to select the next optimal tasks from the graph.
-This ongoing selection process prioritizes activities, keeping both arms continuously engaged, either cooperatively or independently, depending on task requirements.
-By emphasizing the directed nature of these task sequences, the methodology enhances operational efficiency, minimizes idle time, and ensures a smooth workflow, significantly improving the robot’s performance in complex operational environments.
-An overview of the DAG-Plan pipeline is illustrated in Figure 1.
-3.1 Directed Acyclic Sub-task Dependency Graph Generation
+III-A Directed Acyclic Sub-task Dependency Graph Generation
 In prior research, task planning for robots typically involves generating a linear sequence of sub-tasks.
@@
 The arm number of node $n_{i}\in N$ represents the arm number of node needed.
-If $n_{i}=1$ a node requires one arm.
-If $n_{i}=2$, both arms are needed for execution.
 The occupy-release pairs are crucial structures in dual-arm task graphs.
@@
 If the graph contains incomplete grasp-release pairs or is not fully connected, we reflect the LLMs to regenerate the task graph.
-Figure 2: The process of Task Planning Inference.
-In the task “cut the carrot and red bell pepper", DAG-Plan initializes common candidate nodes based on the DAG.
-It evaluates node combinations, checks feasibility, and calculates costs.
-The right arm is selected to grasp the carrot and the left to grasp the red bell pepper.
-After execution, the task graph and nodes are updated, adding subsequent release nodes to the priority candidate nodes for each arm.
-In stage 2, each arms are assigned corresponding priority candidate nodes, checked, and executed.
-The task graph and nodes are updated again.
-The priority nodes become empty, indicating the arms are free.
-In stage 3, with only a single node left, the closest arm to the target executes this node.
-3.2 Task Planning Inference with Generated Directed Acyclic Graph
+III-B Task Planning Inference with Generated Directed Acyclic Graph
 After generating the dual-arm task graph, the planning process enters the inference phase.
@@
 The priority candidate nodes include subsequent nodes when the arm is already engaged in an occupy-release pair.
-The common candidate nodes are initially selected by identifying nodes within the graph that no other nodes point to.
+The common candidate nodes are initially selected by identifying nodes within graph that no other nodes point to.
 During execution, once a node completes its operation, it and its associated edges are removed from the graph.
@@
 This strategic selection and execution framework ensures efficient operation and task handling by the dual-arm robotic system.
+Figure 2: The process of Task Planning Inference.
+In task 2 “clean the table (Hard)”, DAG-Plan initializes common candidate nodes based on the DAG.
+It evaluates node combinations, checks feasibility, and calculates costs.
+The right arm is selected to grasp apple and the left to open drawer.
+After execution, the task graph and nodes are updated, adding subsequent release nodes to the priority candidate nodes for right arm.
+In stage 2, right arm is assigned corresponding priority candidate nodes, checked and left arm still selected node in common candidate nodes with empty priority candidate nodes.
+The task graph and nodes are updated again.
+In stage 3, left arm put mug into drawer and right arm grasp lemon.
 Once we obtain the candidate nodes for each robotic arm, we generate all possible combinations of left and right arm candidate nodes.
@@
 If an occupy node is found, we further examine whether its successor nodes contain dependencies that require other conditions to be met first.
-This could lead to prolonged arm occupancy, thereby decreasing operational efficiency.
-The second check assesses the distance between target objects for the left and right arms.
-If these targets are beyond a specified distance threshold, it becomes difficult for both arms to operate simultaneously.
-The third check evaluates the relative positions of the target locations for each arm.
-If the target position for the left arm is to the right of the right arm’s target, there is a risk of the arms crossing and colliding, which would prevent the execution of these candidate nodes.
+This could lead to prolonged arm occupancy even dead-lock, thereby decreasing operational efficiency.
+The second check evaluates the distance between the target objects assigned to the left and right arms.
+If the distance exceeds a predefined threshold, simultaneous operation of both arms becomes infeasible.
+The third check analyzes the relative positions of the target locations for each arm.
+If the target position for the left arm is located to the right of the right arm’s target or is too far away, the arms may collide or fail to reach their targets.
+To acquire the object positions necessary for decision-making, DAG-Plan utilizes Grounding DINO[31] and SAM2[32] to obtain the point clouds corresponding to the specified objects, thereby determining their locations.
+For specific categories of objects such as bottles and bowls, the category-specific 6D pose estimation model SAR-Net[33] is employed to estimate their poses.
 After completing the checks, we calculate the cost of the left and right hand candidate nodes based on the environment state.
-We aim for the target objects to be close to the robotic arms and to each other, facilitating dual-arm operations.
-Therefore, the cost is composed of two parts: the distance from each target to the respective arm, and the distance between the targets.
+We aim for the target objects to be close to the robotic hands facilitating dual-arm operations.
 The pair with the lowest combined cost is selected for execution.
 The cost $J$ is represented as:
-$J=\text{dis}(obj_{\text{right}},obj_{\text{left}})+\frac{1}{2}\left[\text{dis}(obj_{\text{right}},hand_{\text{right}})+\text{dis}(obj_{\text{left}},hand_{\text{left}})\right].$
-3.3 Motion Planning and Reinforcement Learning Mixed Skill Learning
+$\footnotesize J=\text{dis}(obj_{\text{right}},hand_{\text{right}})+\text{dis}(obj_{\text{left}},hand_{\text{left}}).$
+III-C Sub-tasks Execution with Foundation Model
 Once the sub-tasks to be executed are determined, the robot needs to perform the corresponding actions to bridge the gap between textual instructions and the physical environment.
-Reinforcement Learning (RL) is a powerful technique that enables robots to acquire and refine skills autonomously by interacting with their environment[42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54].
-By engaging in a trial-and-error process and receiving feedback in the form of rewards, robots can learn to optimize their actions to achieve specific objectives [55, 56, 57, 58, 59, 60].
-Recent advancements [61, 62, 63, 64] have further enhanced automation by using LLMs to generate text-to-reward mappings.
-We combine motion planning (RRT-connect) [65] and RL (PPO) [47] to facilitate skill learning (see details in Appendix A).
-The robot approaches the target object using motion planning and then learns the corresponding skill through RL.
-We input the generated dual-arm task graph and description of the reward APIs into LLMs, which then combine different reward APIs and determine if they are necessary conditions for completing the task, thus achieving skill learning.
-However, given the sensitivity of RL training and the ambiguity of language, RL policies may fail to achieve the goal or achieve it in unexpected ways.
-Therefore, we manually fine-tune the generated rewards to ensure that the final skills meet the task requirements.
-4 Experiments
-4.1 Experimental Setup
+The collective perceptual inferences specific to objects, along with physical insights and parameters for manipulation, are methodically organized and converted into a structured format of executable action code.
+III-C1 Occupy
+We use the pre-trained AnyGrasp[20] to produce a variety of grasp pose proposals.
+III-C2 Tool use
+We have specifically designed tool-usage skills based on the categories of tools employed, and have provided adjustable parameters to enable the robot to perform flexible tool-manipulating actions.
+III-C3 Release
+Based on the current grasping position of the object and the 3D bounding box of the target object, we propose an appropriate release pose to ensure that the target object can be accurately placed at the desired location.
+III-C4 Operate
+We employ the GAPartNet[21] to forecast the physical characteristics of articulated objects.
+This model works by dividing the point cloud of an articulated object into its constituent rigid parts and subsequently calculating the articulation parameters.
+Figure 3: Snapshots of 5 Tasks of Dual-arm Kitchen Benchmark.
+TABLE I: Task list of Dual-arm Kitchen Benchmark.
+Index Task Name Human Instruction
+Task 1 Clean the table (Easy) Clean the table. Put objects into plate.
+Task 2 Clean the table (Hard) Clean the table. The fruits should into the plate. The table should be wiped by sponge. The mug and sponge should in drawer.
+Task 3 Stack bowls Stack the bowls onto the wooden tray with the green bowl, blue bowl and yellow bowl order.
+Task 4 Make cup of coffee Make a cup of coffee. You should add the coffee, water and milk in order. Finally, stir it with the spoon.
+Task 5 Boil vegetables Boil vegetables. Pour water and put vegetables into the pot.
+IV Experiments
+In the experiments, we are interested in answering the following questions about DAG-Plan:
+  • $\bullet$
+    To what extent can the task-graph approach improve performance compared to the task-sequence method?
+  • $\bullet$
+    How does the cost of DAG-Plan compare to that of iterative methods?
+  • $\bullet$
+    How does the performance of DAG-Plan compare to multi-agents methods?
+  • $\bullet$
+    Can DAG-Plan be deployed in real-world setting?
+IV-A Experimental Setup
 To validate the correctness and execution efficiency of our method, we created a Dual-arm Kitchen Benchmark, including plan tests and physical simulation tests.
-This benchmark fills the gap in long-sequence operation rigid body physics simulation benchmarks for dual-arm robots.
-The goal of this benchmark is to validate the success rate and efficiency of dual-arm robot planning in complex scenarios.
+This benchmark fills the gap in long-sequence operation physics simulation benchmarks for dual-arm robots.
+The goal of this benchmark is to validate the success rate and efficiency of dual-arm robot planning and executable ability in complex scenarios.
 The dual-arm robot should correctly plan and fully utilize both the left and right arms, completing tasks with as few execution stages as possible.
-The benchmark consists of 9 sequential tasks (see details in Appendix B) are shown in Table 1, comprising a total of 78 sub-tasks.
-The main setting is a modern kitchen divided into three scenarios shown in Figure 3, containing a total of 26 assets and objects.
+The benchmark consists of 5 sequential tasks are shown in Table I and Figure 3, comprising a total of 44 sub-tasks.
 Our physical simulation scene is built on the Sapien [34].
-The embodied platform for planning execution is a dual-arm robot with a mobile base, each 7-degree-of-freedom arm equipped with a two-finger gripper.
-The platform is equipped with five RGB-D cameras, attached above and below each arm’s gripper and on the robot’s head.
-Table 1: Task list of Dual-arm Kitchen Benchmark.
-Index Instruction Index Instruction
-Task 1 Put the apple and bread onto the plate Task 6 Heat the soup and pour a cup of cola
-Task 2 Place the apple on the plate and toast the bread Task 7 Put the apple and the pear into the bowl
-Task 3 Juice the apple and toast the bread Task 8 Cut the carrot and red bell pepper
-Task 4 Wash the cup and the bowl Task 9 Make a pot of soup
-Task 5 Heat the soup and put the tin on the table
-Figure 3: Snapshots of 3 Scenarios of Dual-arm Kitchen Benchmark. Scenario 1 encompasses tasks 1-3, scenario 2 encompasses tasks 4-6, and scenario 3 encompasses tasks 7-9.
-Evaluation of Planning Effectiveness and Conciseness.
+The embodied platform is Agilex CobotMagic, each 6 DOF arm equipped with a two-finger gripper.
+The platform is equipped with an Intel RealSense L515 RGB-D camera, attached on the robot’s head.
+IV-A1 Evaluation of Planning Effectiveness and Conciseness
 In this experiment, we focus on testing the conciseness of the generated plans and the number of stages required.
@@
 This requires that the plans generated by the LLMs can achieve the task goals in terms of language logic and do not violate the preconditions for stage execution.
-Since the lower-level execution method uses reinforcement learning, learning from incorrect plans is costly and unproductive, making plan validation crucial.
-We used GPT-4111https://openai.com/api.
-This work mainly uses gpt-4-turbo-2024-04-09.
-under the temperature of 0.5 to generate five plans for each task, evaluating their Success Rate (SR) and the fewest number of Stage of the passed plan required at the language level.
+We used LLMs to generate 10 plans for each task, evaluating their Success Rate (SR) and the minimum Stage of the passed plan required at language level.
+Due to cost considerations, we also include average Tokens usage as an important metric for evaluating cost of the algorithm.
 Finally, we calculated the average success rate and the average number of stages for all tasks.
@@
 For failed dual-arm plans, we calculated the stage count based on the single-arm plan stages to ensure a fair comparison.
-Evaluation with Physical Simulation.
+IV-A2 Evaluation with Physical Simulation
 In this experiment, we will test the executability and execution efficiency of the plans in physical simulation scenarios.
 Compared to plan tests, physical simulation tests validate both the high-level planning and low-level execution capabilities.
-We will use reinforcement learning to acquire executable skills from the plans that passed the plan tests and then execute them in the physical simulation environment.
-We select the plan that passes the plan test and requires the minimum stages for skill learning.
-Then evaluate the Success Rate (SR) of execution and minimum execution Time in the physical environment with 10 trials.
-Finally, we calculated the average success rate and the execution time for all tasks.
-We defined Execution Efficiency as the ratio of single-arm plan execution time to the execution time required by each method.
-For failed dual-arm plans, we calculated the average time based on the single-arm execution time to ensure a fair comparison.
-Task Planning for Single-arm (TP-S) directly uses LLMs to generate a full task list, with each stage involving a single arm or both arms to manipulate a single object.
-Task Planning for Dual-arm (TP-D) also directly uses LLMs to generate a full task list, but each stage can use the arms to manipulate either a single object or two different objects.
-Our method, DAG-Plan, generates a task graph, followed by task planning inference to iteratively generate nodes for each stage.
-4.2 Experimental Results
-Evaluation of Planning Effectiveness and Conciseness.
-As shown in Table 2, in the plan tests, DAG-Plan consistently outperformed both TP-S and TP-D, showcasing superior efficiency and robustness.
+Then evaluate the Success Rate (SR) and minimum Time which consists of query time and execution time.
+in the physical environment with 10 trials.
+Finally, we calculated the average success rate and the time for all tasks.
+We defined Execution Efficiency as the ratio of single-arm plan time to the time required by each method.
+For failed dual-arm plans, we calculated the average time based on the single-arm time to ensure a fair comparison.
+IV-A3 Baseline Algorithms and Proposed Method
+We compare the planning algorithms baselines and our proposed method:
+  1. 1.
+    Task Planning for Single-arm (TP-S): TP-S directly uses LLMs to generate a full task sequence, with each stage involving a single arm to manipulate a single object.
+  2. 2.
+    Task Planning for Dual-arm (TP-D): TP-D directly uses LLMs to generate a full task sequence, with each stage involving dual arms.
+  3. 3.
+    Twostep [11]: Twostep decomposes complex tasks into fully independent sub-tasks with dual agents, with the primary agent and the secondary agent each completing separate sub-tasks to accelerate task execution.
+  4. 4.
+    RoCo [12]: RoCo proposed multi-agents collaboration using LLMs for planning which is an iterative method.
+Robots query LLMs in each stages to collaboratively reason about task strategies.
+DABICO [13] is a dual-armed version of the RoCo, and is equivalent to the RoCo in our experiments settings.
+  5. 5.
+    DAG-Plan (Proposed): Our method DAG-Plan generates a task graph, followed by task planning inference to iteratively generate nodes for each stage.
+IV-B Evaluation of Planning Effectiveness and Conciseness
+Figure 4: Simulation snapshots of the execution process of long-horizon task 2.
+TABLE II: Performance comparison on plan tests. We report the success rate, minimum stage and average token usage of the 10 plans generated by the LLM for each task.
+Task1 Task2 Task3 Task4 Task5 Average
+[…]
+TP-S $100\%$ $8$ $2030.0$ $70\%$ $11$ $2424.8$ $100\%$ $6$ $2086.8$ $90\%$ $12$ $2413.0$ $100\%$ $7$ $2096.6$ $92\%$ $100.0\%$ $2210.2$
+[…]
+TwoStep[11] $100\%$ $4$ $3699.6$ $0\%$ $Fail$ $4248.2$ $0\%$ $Fail$ $3823.4$ $0\%$ $Fail$ $4130.2$ $100\%$ $5$ $3810.2$ $40\%$ $115.8\%$ $3942.3$
+RoCo[12] $100\%$ $4$ $26615.0$ $0\%$ $Fail$ $68921.4$ $100\%$ $4$ $25410.2$ $40\%$ $8$ $65242.0$ $100\%$ $4$ $26985.0$ $68\%$ $141.9\%$ $42634.7$
+DAG-Plan $100\%$ $4$ $2031.0$ $70\%$ $7$ $2207.2$ $100\%$ $4$ $1980.6$ $70\%$ $7$ $2277.0$ $100\%$ $4$ $1973.0$ $88\%$ $169.2\%$ $2093.8$
+TABLE III: Performance comparison on physical simulation tests. We report the success rate and minimum time for each task with 10 trials. Time consists of query time and execution time.
+Task1 Task2 Task3 Task4 Task5 Average
+[…]
+TP-S $80\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}9.5}+{\color[rgb]{0.78515625,0.11328125,0.19140625}53.5}=63.0$ $40\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}12.9}+{\color[rgb]{0.78515625,0.11328125,0.19140625}90.7}=103.6$ $90\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}8.3}+{\color[rgb]{0.78515625,0.11328125,0.19140625}35.6}=43.9$ $80\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}8.2}+{\color[rgb]{0.78515625,0.11328125,0.19140625}116.9}=125.1$ $100\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}8.2}+{\color[rgb]{0.78515625,0.11328125,0.19140625}55.8}=64.0$ $78\%$ $100.0\%$
+TP-D $80\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}8.5}+{\color[rgb]{0.78515625,0.11328125,0.19140625}26.1}=34.6$ $0\%$ $Fail$ $0\%$ $Fail$ $0\%$ $Fail$ $50\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}8.3}+{\color[rgb]{0.78515625,0.11328125,0.19140625}43.2}=51.4$ $26\%$ $111.4\%$
+TwoStep[11] $80\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}11.4}+{\color[rgb]{0.78515625,0.11328125,0.19140625}27.3}=38.7$ $0\%$ $Fail$ $0\%$ $Fail$ $0\%$ $Fail$ $50\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}11.3}+{\color[rgb]{0.78515625,0.11328125,0.19140625}43.2}=54.5$ $26\%$ $109.2\%$
+RoCo[12] $80\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}49.3}+{\color[rgb]{0.78515625,0.11328125,0.19140625}26.2}=75.5$ $0\%$ $Fail$ $90\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}54.9}+{\color[rgb]{0.78515625,0.11328125,0.19140625}24.1}=79.0$ $30\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}155.3}+{\color[rgb]{0.78515625,0.11328125,0.19140625}78.9}=234.2$ $100\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}54.4}+{\color[rgb]{0.78515625,0.11328125,0.19140625}35.3}=89.7$ $60\%$ $68.7\%$
+DAG-Plan $80\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}8.5}+{\color[rgb]{0.78515625,0.11328125,0.19140625}27.6}=36.1$ $50\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}11.7}+{\color[rgb]{0.78515625,0.11328125,0.19140625}59.8}=71.5$ $90\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}8.3}+{\color[rgb]{0.78515625,0.11328125,0.19140625}24.0}=32.3$ $50\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}8.3}+{\color[rgb]{0.78515625,0.11328125,0.19140625}69.9}=77.9$ $100\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}8.3}+{\color[rgb]{0.78515625,0.11328125,0.19140625}35.4}=43.7$ $74\%$ $152.8\%$
+As shown in Table II, in the plan tests, DAG-Plan consistently outperformed baselines, showcasing superior efficiency and robustness.
 DAG-Plan achieved a high success rate across all tasks, demonstrating its effectiveness in dual-arm manipulation.
-Notably, it maintained an impressive macro average success rate of 97.8% and exhibited a significant reduction in the required stages for task completion, averaging 5.67 stages.
-Furthermore, the completeness of the DAG generation is high, with only an incomplete DAG generated in task 8.
-By re-generating it through reflection, a complete DAG was obtained, increasing the success rate from 0.8 to 1.0.
+Notably, it maintained an impressive macro average success rate of 88% and exhibited a significant reduction in the required stages for task completion, achieving 169.2% stage efficiency.
 These results underscore the effectiveness of DAG-Plan in achieving task goals and its efficiency in execution.
@@
 Although TP-S maintained a relatively high and consistent success rate, it was less efficient in stage minimization.
-Moreover, TP-D, relying on language models to generate dual-arm task plans, exhibited a significantly lower success rate (macro average SR of 44.5%), and often produced plans that were not executable in the physical environment.
+Moreover, TP-D, relying on language models to generate dual-arm task plans, exhibited a significantly lower success rate (macro average SR of 40%), and often produced plans that were not executable in the physical environment.
 While theoretically capable of reducing the number of stages required for tasks, TP-D frequently encountered challenges related to coordination complexities and unrealistic task assignments.
 This highlights the superiority of DAG-Plan in effectively translating high-level plans into physical actions and navigating complexities in the environment with relative ease compared to TP-D.
-Additionally, we provide detailed plans of DAG-Plan and the baseline in Appendix C.
-Table 2: Performance comparison on plan tests. We report the success rate and minimum stage of the 5 plans generated by the LLM for each task, and the macro average metrics for all tasks.
-Scenario1 Scenario2
-Task1 Task2 Task3 Task4 Task5 Task6
-SR Stage SR Stage SR Stage SR Stage SR Stage SR Stage
-TP-S $1.0$ $4$ $0.8$ $6$ $1.0$ $8$ $1.0$ $8$ $1.0$ $10$ $0.8$ $11$
-TP-D $1.0$ $2$ $0.0$ $Fail$ $1.0$ $4$ $0.0$ $Fail$ $0.2$ $6$ $0.6$ $6$
-DAG-Plan $1.0$ $2$ $1.0$ $4$ $1.0$ $4$ $0.8$ $6$ $1.0$ $6$ $1.0$ $6$
-Scenario3 Macro Avg
-Task7 Task8 Task9 SR Stage Stage Efficiency
-SR Stage SR Stage SR Stage
-TP-S $1.0$ $8$ $0.6$ $10$ $1.0$ $13$ $91.1\%$ $8.67\pm 2.54$ $100.0\%$
-TP-D $0.6$ $5$ $0.6$ $7$ $0.0$ $Fail$ $44.5\%$ $6.33\pm 2.87$ $137.0\%$
-DAG-Plan $1.0$ $5$ $1.0$ $7$ $1.0$ $11$ $97.8\%$ $5.67\pm 2.35$ $152.9\%$
-Table 3: Performance comparison on physical simulation tests. We report the success rate and minimum time for each task with 10 trials, and the macro average metrics for all tasks.
-Scenario1 Scenario2
-Task1 Task2 Task3 Task4 Task5 Task6
-SR Time SR Time SR Time SR Time SR Time SR Time
-TP-S $0.7$ $37.1$ $0.7$ $59.0$ $0.7$ $79.6$ $1.0$ $84.1$ $0.3$ $105.5$ $0.3$ $114.4$
-TP-D $1.0$ $18.6$ $0.0$ $Fail$ $0.6$ $40.1$ $0.0$ $Fail$ $0.0$ $Fail$ $0.0$ $Fail$
-DAG-Plan $1.0$ $18.6$ $0.7$ $39.4$ $0.6$ $40.1$ $1.0$ $66.3$ $0.3$ $63.2$ $0.3$ $76.3$
-Scenario3 Macro Avg
-Task7 Task8 Task9 SR Time Execution Efficiency
-SR Time SR Time SR Time
-TP-S $0.5$ $74.0$ $0.1$ $105.2$ $0.2$ $136.2$ $49.9\%$ $88.3\pm 28.5$ $100.0\%$
-TP-D $0.5$ $55.1$ $0.3$ $76.4$ $0.0$ $Fail$ $26.7\%$ $76.6\pm 35.5$ $115.3\%$
-DAG-Plan $0.6$ $53.3$ $0.3$ $76.4$ $0.2$ $107.4$ $55.6\%$ $60.1\pm 24.5$ $147.0\%$
-Evaluation with Physical Simulation.
-As shown in Table 3, the physical simulation tests provided further insights into the practical applicability and execution capabilities of our planning methods under more dynamic and realistic conditions.
-Here again, DAG-Plan demonstrated a balanced performance with a solid success rate and efficient execution times.
+The key idea of Twostep is to decompose a task into two completely independent sub-tasks.
+While this approach is highly effective in multi-agents task settings, it falls short in dual-arm tasks.
+In dual-arm scenarios, the entire task sequence is interconnected, especially in tasks with order requirements, such as Task 3 and Task 4.
+Each sub-task is inherently dependent on others, making it impossible to decompose the task into fully independent sub-tasks.
+As a result, Twostep achieves low success rates and is ineffective in dual-arm tasks.
+RoCo is capable of acquiring the current environment and robot state at each step.
+Through communication between the left and right arms, it can often select appropriate subtasks, achieving high success rates and phase efficiency for all tasks except Task 2.
+In Task 2, RoCo is still limited by the drawback of sequentially generating dual-arm plans, neglecting implicit dependency conditions.
+For instance, it attempts to grasp the sponge and mug without first opening the drawer, resulting in a failure to open the drawer and ultimately causing the plan to fail.
+Additionally, RoCo requires querying the large model and retaining historical dialogue in each round, leading to significant token consumption in long-sequence tasks.
+The success of DAG-Plan can be attributed to its DAG generation, which systematically breaks down tasks into manageable components.
+This allows for more accurate and feasible plan generation, as the method iteratively refines the task execution steps while maintaining alignment with the operational capabilities of the robotic system.
+The failed case of DAG-Plan is LLMs occasionally generate some unreasonable task graphs.
+For instance, each node depend on the previous node, causing the task graph to degenerate into a task sequence, or certain dependencies be omitted.
+IV-C Evaluation with Physical Simulation
+As shown in Table III, the physical simulation tests provided further insights into the practical applicability and execution capabilities of our planning methods under more dynamic and realistic conditions.
+Here again, DAG-Plan demonstrated a balanced performance with a solid success rate and efficient times.
 We show the execution process of DAG-Plan in a physical simulation environment in Figure 4.
-Moreover, we provide detailed analysis and explanation of DAG-Plan and the baseline in Appendix D.
+Figure 5: Real-world snapshots of the execution process of long-horizon tasks.
 Compared to TP-D, DAG-Plan effectively translates high-level plans into feasible actions based on target object information and the robot’s current state under the guidance of a task graph.
-Both DAG-Plan and TP-S were able to complete $/9$ 9 tasks in the physical simulation tests.
-However, while TP-D passed the plan tests for $6/9$ tasks, it only completed $4/9$ in the physical simulation tests.
-In tasks 5 and 6, the plans generated by TP-D could not be executed due to real-world physical constraints.
-For example, in task 6, the plan of TP-D included “put cola bottle into refrigerator cooler" with left arm and “close microwave door" with right arm.
-However, since the refrigerator is on the right side of the microwave, the dual-arm robot could not cross its arms to execute this action.
-Ultimately, DAG-Plan achieved a 28.9% higher success rate in physical simulation tests compared to TP-D.
+Both DAG-Plan and TP-S were able to complete $/5$ 5 tasks in the physical simulation tests.
+However, while TP-D passed the plan tests for $3/5$ tasks, it only completed $2/5$ in the physical simulation tests.
+In tasks 5, the plans generated by TP-D could not be executed due to real-world physical constraints.
+In task 5, the plan of TP-D included “grasp red bell pepper” with left arm and “grasp carrot” with right arm.
+However, since the red bell pepper and carrot both on the table’s right side and too close, the left-arm could not grasp red bell pepper.
+Instead DAG-Plan select sub-tasks “grasp water bottle” with left arm and “grasp carrot” with right arm.
+Ultimately, DAG-Plan achieved a 48% higher success rate in physical simulation tests compared to TP-D.
 This demonstrates that, both in high-level planning and low-level execution, the success rate of the dual-arm plans generated by TP-D is significantly inferior to those generated by DAG-Plan.
+Furthermore, although RoCo’s stage efficiency is high, it is limited by the fact that each stage needs to be associated with query LLMs, which makes its query time much higher than that of the non-iterative approach, resulting in inefficiencies in actual execution.
 Regarding execution efficiency, the dual-arm plan allows for parallel execution of sub-tasks, resulting in higher efficiency.
-DAG-Plan’s execution efficiency was 47.0% higher than TP-S, as it maximized the parallelization of sub-tasks while ensuring the feasibility of the plan.
-In all tasks, DAG-Plan’s execution time was shorter than TP-S.
-Although TP-D had similar execution times to DAG-Plan for some tasks, its overall efficiency was low because $5/9$ tasks could not be executed.
-Consequently, TP-D’s efficiency was only 15.3% higher than TP-S.
-Overall, the experimental results strongly support the adoption of DAG-Plan in complex robotic operations.
-DAG-Plan not only outperforms traditional single-arm and dual-arm planning approaches in terms of success rates and efficiency but also demonstrates significant robustness and reliability in translating plans into actionable steps in a physical context.
-Figure 4: Head camera snapshots of the execution process of 2 example long-horizon tasks.
-5 Conclusion
-This work introduces DAG-Plan, which efficiently and accurately generates collaborative plans with LLMs for mobile dual-arm robots.
+DAG-Plan’s execution efficiency was 52.8% higher than TP-S, as it maximized the parallelization of sub-tasks while ensuring the feasibility of the plan.
+Despite TP-D had similar execution times to DAG-Plan for some tasks, its overall efficiency was low because $3/5$ tasks could not be executed.
+Consequently, TP-D’s efficiency was only 11.4% higher than TP-S.
+Although RoCo has a high success rate and stage efficiency, it is limited by query time which is even longer than execution time, and the execution efficiency is only 68.7%.
+IV-D Real-world Experiments
+We validate DAG-Plan in real-world setting, where a Agilex Cobot Magic dual-arm robot to complete tasks similar to simulation.
+When dealing with these tasks, the DAG-Plan framework can effectively assign the sub-task for the each arm.
+We show the execution process of DAG-Plan in real-world in Figure 5 first row.
+We evaluate 10 runs for each task as shown in Figure 6 for comparison of real-world and simulation success rates.
+The slightly lower success rate in real-world is primarily due to the inaccuracy of depth, whereas the depth in simulations is entirely precise.
+Figure 6: Comparison of real-world and simulation success rates of DAG-Plan.
+In addition to the predefined tasks similar to simulation, we also conducted tests in the wild, as shown in Figure 5 second row.
+By providing human instructions through audio input, the LLM automatically generates environment descriptions and DAG, enabling the dual-arm robot to efficiently complete the tasks.
+V Conclusion
+This work introduces DAG-Plan, which efficiently and accurately generates collaborative plans with LLMs for dual-arm robots.
 DAG-Plan decomposes complex tasks into directed acyclic graph (DAG) with clear temporal relationships and iteratively selects feasible sub-tasks based on environmental observations during execution.
-The main contribution is the replacement of task sequences with a DAG and the dynamic adjustment of the planning according to the current situation, allowing dual-arm robots to flexibly utilize both arms for sub-task execution.
-We also conducted a dual-arm kitchen benchmark, providing a testing scenario for future long-horizon dual-arm works.
-Limitations and Future Works.
-Currently, DAG-Plan relies on reinforcement learning for underlying skills, which may be inefficient for complex tasks.
-Our future efforts will aim to enhance skill learning modules, boosting automation and success rates for real-world applications.
-Appendix A Implementation Details of Motion Planning and Reinforcement Learning
-A.1 Embodied Platform
-We used the Baxter robot from Rethink Robotics as the primary embodied platform shown in Figure 5.
-We configured its base in the simulation as a mobile base capable of translating on the plane but not rotating.
-The grippers of the Baxter robot were replaced with Robotiq 85 grippers to achieve more robust grasping.
-Figure 5: Embodied platform of DAG-Plan.
-A.2 Motion Planning
-For motion planning, we use open-source implementations of RRT-connect [^2] https://github.com/haosulab/mplib.
-We manually set multiple candidate poses and use motion planning to find the pose that requires the shortest movement steps while avoiding collisions.
-For sub-tasks of the types occupy, tool use, and operate, we use motion planning to move the end-effector to within 0.15 cm of the target object, facing it as directly as possible.
-For release type sub-tasks, we adjust based on the opening direction of the target object.
-For high-precision sub-tasks that require both arms, such as “put pot into sink”, where motion planning alone struggles to maintain fixed distance and equal height for stable placement, we directly use reinforcement learning to learn the sub-task.
-A.3 Reinforcement Learning
-For reinforcement learning training, we use open-source implementations of PPO [^3] https://github.com/DLR-RM/stable-baselines3 algorithm, and list the hyper-parameters in Table 4.
-The observation space is the low-level state of the objects and robot and the images captured by 5 RGB-D cameras.
-The control mode of mobile base adopts velocity control.
-The control mode of the dual-arm adopts arm joint delta position control, which means the action space consists of the change of the arm joint position.
-Table 4: Hyper-parameter of PPO algorithm.
-Hyper-parameter Value
-Discount factor $\gamma$ $0.99$
-# of epochs per update $10$
-[…]
-# of environments $32$
-Batch size $400$
-Target KL divergence None
-[…]
-# of steps per update $4800$
-Rollout steps per episode $100$
-Corresponding rollout seconds per episode $4s$
-Shape of observation space $(128,128,20)$Visual input
-$95$State input
-Shape of action space $18$
-In this work, we use a set of reward primitives shown in Table 5.
-We use LLMs to combine reward primitives and specify the corresponding object and target object, with manual adjustments for corrections to generate dense rewards.
-The reward function and success condition for each sub-task are shown in Table 6.
-In the final reward, we normalize the reward functions of the subtasks corresponding to the left and right hands to 1 and assign additional rewards for successful states.
-The reward $r$ is represented as:
-$r=r_{\text{right}}+r_{\text{left}}+r_{\text{side\_success}}\left\{\begin{aligned} &30,\text{if side success first time}\\ &0,\enspace\text{otherwise}\end{aligned}\right.+r_{\text{success}}\left\{\begin{aligned} &150-\text{steps},\text{if success}\\ &0,\quad\quad\quad\quad\enspace\text{otherwise}\end{aligned}\right.$
-Table 5: Reward primitives.
-Reward Primitives Description Pseudocode
-$obj\_dis$ Move object close to target object $\frac{1}{2}(1-\frac{\text{dis}(obj,target)}{init\_dis}+1-\text{tanh}(\text{dis}(obj,target)))$
-[…]
-$obj\_grasped$ Grasp target object $1\enspace if\enspace\text{grasped}(target)\enspace else\enspace 0$
-$obj\_in\_obj$ Put object into target object $1\enspace if\enspace\text{in}(obj,target)\enspace else\enspace 0$
-$obj\_on\_obj$ Put object onto target object $1\enspace if\enspace\text{on}(obj,target)\enspace else\enspace 0$
-$joint\_qpos$ Move joint to target joint position $\frac{\text{clip}(pos_{cur},pos_{min},pos_{max})-pos_{min}}{pos_{max}-pos_{min}}$
-$ee\_pos$ Move end-effector to target object $\frac{1}{2}(1-\frac{\text{dis}(hand,target)}{init\_dis}+1-\text{tanh}(\text{dis}(hand,target)))$
-$ee\_height$ Keep both end-effectors at same height $1-\frac{\text{min}((\text{abs}(z_{right}-z_{left})),0.05)}{0.05}$
-$obj\_cut$ Cut the target object with object (knife) $\frac{1}{2}(obj\_grasped(obj)+obj\_on\_obj(obj,target))$
-$obj\_pour$ Pour liquid in object into target object $\frac{1}{3}(obj\_grasped(obj)+obj\_horizon(obj)+obj\_above(obj,target))$
-Table 6: Reward and success condition of Sub-tasks. Success condition is marked in bold.
-Sub-task Pseudocode
-$grasp<target\_obj>$ $1.0\times ee\_pos(target)+1.0\times\mathbf{obj\_grasped(target)}$
-$pour<obj>into<target\_obj>$ $1.0\times\mathbf{obj\_pour(obj,target)}$
-$cut<target\_obj>with<obj>$ $1.0\times\mathbf{obj\_cut(obj,target)}$
-$put<obj>into<target\_obj>$ $3.0\times obj\_dis(obj,target)+1.0\times\mathbf{obj\_in(obj,target)}$
-$put<pot>into<sink>$ $10.0\times ee\_both(sink)+3.0\times\mathbf{obj\_grasped(pot)}+1.0\times ee\_height+1.0\times\mathbf{obj\_in(pot,sink)}+1.0\times\mathbf{obj\_horizon(pot)}$
-$put<obj>onto<target\_obj>$ $3.0\times obj\_dis(obj,target)+1.0\times\mathbf{obj\_on(obj,target)}$
-$put<pot>onto<stove\_body>$ $10.0\times ee\_both(stove\_body)+3.0\times\mathbf{obj\_grasped(pot)}+1.0\times ee\_height+1.0\times\mathbf{obj\_on(pot,stove\_body)}+1.0\times\mathbf{obj\_horizon(pot)}$
-$open<target\_obj>$ $3.0\times\mathbf{joint\_qpos(target\_joint,pos_{target})}+1.0\times obj\_dis(obj,target)+0.5\times obj\_grasped(obj,target)$
-$close<target\_obj>$ $3.0\times\mathbf{joint\_qpos(target\_joint,pos_{target})}+1.0\times obj\_dis(obj,target)+0.5\times obj\_grasped(obj,target)$
-$switch\enspace on<target\_obj>$ $3.0\times\mathbf{joint\_qpos(target\_joint,pos_{target})}+1.0\times obj\_dis(obj,target)+0.5\times obj\_grasped(obj,target)$
-$switch\enspace off<target\_obj>$ $3.0\times\mathbf{joint\_qpos(target\_joint,pos_{target})}+1.0\times obj\_dis(obj,target)+0.5\times obj\_grasped(obj,target)$
-Appendix B Dual-arm Kitchen Benchmark Details
-B.1 Sub-tasks Action Capabilities
-The action capabilities of sub-tasks in the Dual-arm Kitchen Benchmark are listed in Table 7.
-Table 7: Sub-tasks action capabilities. Different actions correspond to the node types occupy, tool use, release, and operate.
-Sub-task Description
-$grasp<target\_obj>$ Grasp target object
-$pour<obj>into<target\_obj>$ Pour the liquid in object into target object
-$cut<target\_obj>with<obj>$ Cut target object which is food with object which is a knife
-$put<obj>into<target\_obj>$ Put object in hand into target object
-$put<obj>onto<target\_obj>$ Put object in hand onto target object
-$open<target\_obj>$ Open target object, such as a door, a drawer floor, a faucet
-$close<target\_obj>$ Close target object, such as a door, a drawer floor, a faucet
-$switch\enspace on<target\_obj>$ Switch on target object which is an electric device
-$switch\enspace off<target\_obj>$ Switch off target object which is an electric device
-B.2 Full Sub-tasks List
-The sub-tasks of each task in the Dual-arm Kitchen Benchmark are listed in Table 9.
-Table 8: Sub-tasks of Dual-arm Kitchen Benchmark.
-Index Instruction Sub-tasks
-Task 1 Put the apple and bread onto the plate Grasp apple, Put apple onto plate, Grasp bread, Put bread onto plate
-Task 2 Place the apple on the plate and toast the bread Grasp apple, Put apple onto plate, Grasp bread, Put bread into toaster body, Switch on toaster, Switch off toaster
-Task 3 Juice the apple and toast the bread Grasp apple, Put apple into juicer container, Switch on juicer, Switch off juicer, Grasp bread, Put bread into toaster body, Switch on toaster, Switch off toaster
-Task 4 Wash the cup and the bowl Open dishwasher door, Grasp cup, Put cup into dishwasher rack, Grasp bowl, Put bowl into dishwasher rack, Close dishwasher door, Switch on dishwasher, Switch off dishwasher
-Task 5 Heat the soup and put the tin on the table Open refrigerator cooler door, Grasp bowl, Open microwave door, Put bowl into microwave body, Close microwave door, Switch on microwave, Switch off microwave, Grasp tin, Put tin onto table, Close refrigerator cooler door
-Task 6 Heat the soup and pour a cup of cola Open refrigerator cooler door, Grasp bowl, Open microwave door, Put bowl into microwave body, Close microwave door, Switch on microwave, Switch off microwave, Grasp cola bottle, Pour cola bottle into cup, Put cola bottle into refrigerator cooler body, Close refrigerator cooler door
-Task 7 Put the apple and the pear into the bowl Open cabinet left door, Grasp bowl, Put bowl onto table, Grasp apple, Put apple into bowl, Grasp pear, Put pear into bowl
-Task 8 Cut the carrot and red bell pepper Open cabinet left door, Grasp knife, Grasp carrot, Put carrot onto cutting board, Cut carrot with knife, Grasp red bell pepper, Put red bell pepper onto cutting board, Cut red bell pepper, Put knife onto cutting board
-Task 9 Make a pot of soup Open cabinet right door, Grasp pot, Put pot into sink, Open faucet, Close faucet, Grasp pot, Put pot onto stove body, Switch on stove, Grasp carrot, Put carrot into pot, Grasp red bell pepper, Put red bell pepper into pot, Switch off stove
-B.3 Assets of Dual-arm Kitchen Benchmark
-The assets of each task in the Dual-arm Kitchen Benchmark are listed in Table 9.
-These assets were sourced from Mobility-Partnet [^4] https://sapien.ucsd.edu/browse, BlenderKit [^5] https://www.blenderkit.com, and Sketchfab [^6] https://sketchfab.com.
-We have modified them and constructed the corresponding URDF models.
-Table 9: Assets of Dual-arm Kitchen Benchmark.
-Index Instruction Assets
-Task 1 Put the apple and bread onto the plate table, apple, bread, juicer, toaster
-Task 2 Place the apple on the plate and toast the bread table, apple, bread, juicer, toaster
-Task 3 Juice the apple and toast the bread table, apple, bread, juicer, toaster
-Task 4 Wash the cup and the bowl table, dishwasher, microwave, cup, bowl, refrigerator cooler
-Task 5 Heat the soup and put the tin on the table table, dishwasher, microwave, tin, bowl, refrigerator cooler
-Task 6 Heat the soup and pour a cup of cola table, dishwasher, microwave, cola bottle, bowl, refrigerator cooler
-Task 7 Put the apple and the pear into the bowl table, sink, faucet, stove, cutting board, cabinet, shelf, knife, pot, apple, pear
-Task 8 Cut the carrot and red bell pepper table, sink, faucet, stove, cutting board, cabinet, shelf, knife, pot, carrot, red bell pepper
-Task 9 Make a pot of soup table, sink, faucet, stove, cutting board, cabinet, shelf, knife, pot, carrot, red bell pepper
-Appendix C Planning Evaluation Analysis
-To further illustrate the differences between our approach and the baselines, we analyzed the planning evaluation of TP-S, TP-D, and DAG-Plan for tasks 5 and 8.
-We validated the generated plans using Planning Domain Definition Language (PDDL) written by robotics experts for the corresponding tasks.
-The plans generated by TP-S and TP-D were executed sequentially in the PDDL environment.
-For DAG-Plan, due to the lack of environmental information in this evaluation, only the first check was conducted, and the first candidate node from the candidate list was selected for execution.
-C.1 Planning Evaluation of Task 5
-We provide the successful plan for each method in Figure 6 and the failed plan for TP-D in Figure 7.
-We further analyzed the failed plans of TP-D.
-TP-D made an error, incorrectly using the left hand, which was holding the tin, to close the refrigerator cooler door.
-In addition, TP-D is missing the sub-task of closing the microwave door.
-C.2 Planning Evaluation of Task 8
-We provide the successful plan for each method in Figure 8 and the failed plan for each method in Figure 9.
-We further analyzed the failed plans of each method.
-TP-S did not follow the rules in this task, failing to specify which hand to use for using the knife.
-TP-D made an error due to the complexity of dual-arm tasks, incorrectly using the left hand, which was not holding the knife, to cut.
-DAG-Plan did not follow the rules initially, generating an incomplete DAG, but after reflecting to LLMs for correction, it ultimately produced the same DAG as shown in Figure 8.
-C.3 Failed Plan Analysis of TP-D
-We further analyzed the failed plans of TP-D on the Dual-arm Kitchen Benchmark, as shown in Figure 10.
-It is evident that directly generating sequence dual-arm plans using LLMs often leads to errors, such as the robot holding an object in its hand while attempting to manipulate another object.
-This type of sequence temporal dependency works well for generating single-arm plans but struggles with the complexity of dual-arm planning, making it difficult to produce correct dual-arm plans.
-In contrast, DAG-Plan imposes strict constraints with occupy-release pair can be performed once the robot holds an object.
-This ensures that such errors are prevented, maintaining the conciseness of the plan.
-Figure 6: Successful plan for each method of task 5.
-Figure 7: Failed plan for each method of task 5. Errors are highlighted in red boxes.
-Figure 8: Successful plan for each method of task 8.
-Figure 9: Failed plan for each method of task 8. Errors are highlighted in red boxes.
-Figure 10: Failed plan for TP-D of task 2, task 4, and task 9. Errors are highlighted in red boxes.
-Appendix D Execution Process Analysis
-D.1 Execution Process for DAG-Plan of Task 5
-We provide the full execution process for DAG-Plan of task 5 in Figure 16.
-(a) Stage 1: Open microwave door (left) and Open refrigerator cooler door (right)
-(a) Stage 2: Grasp tin (left) and Grasp bowl (right)
-(a) Stage 3: Put tin onto table (left) and Put bowl into microwave body (right)
-(a) Stage 4: Close microwave door (left) and Close refrigerator cooler door (right)
-(a) Stage 5: Switch on microwave (left)
-(a) Stage 6: Switch off microwave (left)
-Figure 16: Execution process for DAG-Plan of task 5.
-D.2 Execution Process for DAG-Plan of Task 8
-We provide the full execution process for DAG-Plan of task 5 in Figure 23.
-(a) Stage 1: Grasp red bell pepper (left) and Grasp carrot (right)
-(a) Stage 2: Put red bell pepper onto cutting board (left) and Put carrot onto cutting board (right)
-(a) Stage 3: Open cabinet left door (right)
-(a) Stage 4: Grasp knife (right)
-(a) Stage 5: Cut carrot with knife (right)
-(a) Stage 6: Cut red bell pepper with knife (right)
-(a) Stage 7: Put knife onto cutting board (right)
-Figure 23: Execution process for DAG-Plan of task 8.
-D.3 Failed Execution Analysis of TP-D
-Since TP-D generates dual-arm plans without the ability to adjust hand movements in real-time based on environmental information, some plans that pass PDDL validation still fail in the physical simulation environment.
-For instance, as shown in Figure 24, "Open refrigerator cooler door (left)" and "Open microwave door (right)" cannot be executed due to the hands crossing each other’s paths.
-Figure 24: Failed Execution Process of TP-D.
-Appendix E Full Prompts
-E.1 Our Method: Prompt for DAG-Plan
-{mdframed}
-[backgroundcolor=azure,linewidth=0pt] You are a helpful assistant to plan a definite task into Directed Acyclic Graph (DAG) for a dual-arm robot.
-I will give you the initial environment description:
-task: …
-obj_names: […]
-obj_description: …
-You must follow the following criteria:
-1) Create a DAG based on the task name, considering the task complete when all nodes are executed. The task can be complex but should be broken into simple, reasonable nodes.
-2) The node name must be selected in command API and follow the description.
-3) When moving an object, confirm it clearly specify its destination, e.g., ‘put "obj_name" into/onto "obj_name"’
-4) Synthesize the ’edge’ of a node, a list of node’s indices the precondition of this node.
-When generating edges, please match the current node with all other nodes to determine if there are any dependencies.
-5) Your generated edges are not only the preconditions for the node to be executed, but also preconditions to complete the task objectives.
-For example, in task: fill cup with coffee, node: switch on coffee machine is a executed precondition for node: switch off coffee machine to execute, and additionally, node: put cup into coffee machine is also a task precondition for node: switch off coffee machine.
-If this condition is missing, switch off coffee machine may executed before put cup into coffee machine, the container will not filled with coffee, and the task cannot be completed.
-6) Each node has a node type.
-"occupy" means the hand remains occupied after the node such as node: grasp "apple", and "release" means the object is released after this node such as node: put "apple" into "drawer", and "operate" means robot manipulate joint such as node: open "drawer", and "tool use" means robot utilize obj which handling in hand such as node: cut "apple" with "knife".
-7) Each occupy node must have a corresponding release node which means the obj must holding in hand before put obj.
-8) When two hands are needed to grasp or put an object, use a two-handed node directly instead of two one-handed nodes.
-For example, use node:grasp pot instead node: grasp pot_left_handle and node: grasp pot_right_handle.
-9) Ensure the number of arms used matches the "arm_num" term. The "occupy" type nodes in the edge list should not exceed two.
-10) Enclose all object references in quotes "". These must be in obj_names. Do not introduce new objects or remove any.
-11) You need to generate as few nodes as possible in order to complete the task as quickly as possible while still being able to reach the end goal.
-Any node that is not relevant to the completion of the task should not be present.
-12) After the graph is generated, create a task complete node.
-List of commands of the API, all the object must selected in obj_names:
-* The following command allows robot to open an <target-object> with a door, a drawer floor, a faucet:
-  command syntax:
-  -open <target-object>
-  -close <target-object>
-  node type: operate
-* The following command allows robot to switch on or off an electric device.
-If the electric device has a door, such as a mircrowave with door, this command can only be selected when the door of electric device is closed:
-  command syntax:
-  - switch_on <target-object>
-{mdframed}[backgroundcolor=azure, linewidth=0pt]   - switch_off <target-object>
-  node type: operate
-* The following command allows robot to put an <object> in or on a <target-object>, the <object> must holding in hand, the <target-object> should not holding in hand:
-  command syntax:
-  - put <object> into <target-object>
-  - put <object> onto <target-object>
-  node type: release
-* The following command allows robot to pour a liquid from an <object> into an <target-object>. This command must be used to pour bottle of liquid:
-  command syntax:
-  - pour <object> into <target_object>
-  node type: tool use
-* The following command allows robot to cut a food which is <target-object> with an <object>.
-This <target-object> must on the cutting_board, after cutting the <target-object> keep same name before cutting, and still on cutting board not in hand:
-  command syntax:
-  -cut <target-object> with <object>
-  node type: tool use
-* The following command allows robot to take an <target-object>:
-  command syntax:
-  grasp <target-object>
-  node type: occupy
-Here’s an example input and response:
-INPUT:
-task: put apple into drawer
-obj_names:  ["apple", "drawer_shell","drawer_floor", "table"]
-obj_description:
-    "apple":
-        type: actor
-        fixed: False
-        spatial_relationship: "apple" is on "table".
-        description: None
-    "drawer":
-        type: articulation
-        fixed: True
-        links: ["drawer_shell","drawer_floor","drawer_handle"]
-        joints: ["drawer_joint"]
-        spatial_relationship: "drawer" is on "table". "drawer" is closed now.
-        description:
-        "drawer_shell": This is the external part of the drawer, which cannot be manipulated but can be used to place items.
-The drawer shell is typically the external frame or casing of the drawer, providing structural support.
-        "drawer_floor": This is the internal part of the drawer, containing a "drawer_handle" and "drawer_joint." The drawer floor can be opened or closed by manipulating the "drawer_handle."
-         "drawer_handle": This is a handle located inside the drawer, and by operating it, the drawer can be opened or closed.
-Typically, the drawer handle is a component connected to the internal mechanism of the drawer.
-        "drawer_joint": The parent link of this joint is "drawer_shell".
-The child link of this joint is "drawer_floor".
-It allows movement of the "drawer_floor" relative to the "drawer_shell," enabling the functionality of opening and closing the drawer.
-{mdframed}[backgroundcolor=azure, linewidth=0pt]     "table":
-        type: actor
-        fixed: True
-        spatial_relationship: "table" on the ground.
-        description: None
-RESPONSE:
-nodes:
-    node_1:
-        name: open "drawer_floor"
-        arm_num: 1
-        edge: []
-        type: operate
-        description: Open the drawer to access inside.
-    node_2:
-        name: grasp "apple"
-        arm_num: 1
-        edge: []
-        type: occupy
-        description: Grasp the apple on the table.
-    node_3:
-        name: put "apple" into "drawer_floor"
-        arm_num: 1
-        edge: [1,2]
-        type: release
-        description: Put the apple into the drawer.
-    node_4:
-        name: task complete
-        arm_num: 0
-        edge: [3]
-        type: complete
-        description: Task complete after node_3 has been executed.
-Now, let’s begin:
-E.2 Baseline: Prompt for TP-S
-{mdframed}
-[backgroundcolor=azure, linewidth=0pt] You are a helpful assistant to plan a definite task into multi-satge for a dual-arm robot.
-To be on accuracy side, you can only use single arm or manipulate same object use dual arm each stage.
-I will give you the initial environment description:
-task: …
-obj_names: […]
-obj_description: …
-You must follow the following criteria:
-1) Break down the task into stages. The task can be complex but should be broken into simple, reasonable stages. The (side) should never be omitted.
-2) The stage name must adhere the format: API (side). The API must select in command API and follow the description.
-3) When moving an object, confirm it is grasped by hand and clearly specify its destination, e.g., ’put "obj_name" into/onto "obj_name"’.
-4) Synethesis current hand empty or holding obj after stage has been executed. For example, right_hand: holding "apple", left_hand: empty.
-{mdframed}[backgroundcolor=azure, linewidth=0pt] 5) The next stage you generate should accord to hand.
-For example, right_hand: holding "apple", left_hand: empty.
-The stage_next: put "apple" into "drawer", is reasonable then your right hand is empty, you can use your right_hand.
-The stage_next: grasp "pear" is wrong because your right hand state is holding "apple".
-6) Each grasp stage should have a corresponding put stage which means a stage: grasp obj must have a corresponding stage: put obj.
-7) When the obj in your hand has been used up, it should be returned to a fitting position to release your hands.
-8) When two hands are needed to grasp or put an object, use a two-handed stage directly instead of two one-handed stages.
-For example, use stage: grasp pot (both) instead stage: grasp pot_left_handle (left) and stage_next: grasp pot_right_handle (right).
-9) Enclose all noun references in quotes "". These must be in obj_names or ee_names. Do not introduce new objects or remove any.
-10) You need to generate as few stages as possible in order to complete the task as quickly as possible while still being able to reach the end goal.
-Any stage that is not relevant to the completion of the task should not be present.
-List of commands of the API, all the object must selected in obj_names:
-* The following command allows robot to open an <target-object> with a door, a drawer floor, a faucet:
-  command syntax:
-  -open <target-object>
-  -close <target-object>
-* The following command allows robot to switch on or off an electric device.
-If the electric device has a door, such as a mircrowave with door, this command can only be selected when the door of electric device is closed:
-  command syntax:
-  - switch_on <target-object>
-  - switch_off <target-object>
-* The following command allows robot to put an <object> in or on a <target-object>, the <object> must holding in hand, the <target-object> should not holding in hand:
-  command syntax:
-  - put <object> into <target-object>
-  - put <object> onto <target-object>
-* The following command allows robot to pour a liquid from an <object> into an <target-object>. This command must be used to pour bottle of liquid:
-  command syntax:
-  - pour <object> into <target_object>
-* The following command allows robot to cut a food which is <target-object> with an <object>.
-This <target-object> must on the cutting_board, after cutting the <target-object> keep same name before cutting, and still on cutting board not in hand:
-  command syntax:
-  -cut <target-object> with <object>
-You should only respond in the format as described below and do not respond anything else:
-RESPONSE FORMAT:
-stages:
-[…]
-        name: The name of the stage.
-        right_hand_state: Empty or holding sth.
-        left_hand_state: Empty or holding sth.
-{mdframed}[backgroundcolor=azure, linewidth=0pt] Here’s an example input and response:
-INPUT:
-task: put "apple" into "drawer_floor"
-obj_names:  ["apple", "drawer_shell", "drawer_floor", "table"]
-obj_description:
-    "apple":
-        type: actor
-        fixed: False
-        spatial_relationship: The "apple" is on the "table". The "apple" is close to "right_hand".
-        description: None
-    "drawer":
-        type: articulation
-        fixed: True
-        links: ["drawer_shell","drawer_floor","drawer_handle"]
-        joints:  ["drawer_joint"]
-      spatial_relationship: This is a fixed articulation. The "drawer" is on the "table". The "drawer" is closed now. The "drawer" is close to "left_hand".
-        description:
-        "drawer_shell": This is the external part of the drawer, which cannot be manipulated but can be used to place items.
-The drawer shell is typically the external frame or casing of the drawer, providing structural support.
-        "drawer_floor": This is the internal part of the drawer, containing a "drawer_handle" and "drawer_joint." The drawer floor can be opened or closed by manipulating the "drawer_handle."
-        "drawer_handle": This is a handle located inside the drawer, and by operating it, the drawer can be opened or closed.
-Typically, the drawer handle is a component connected to the internal mechanism of the drawer.
-        "drawer_joint": The parent link of this joint is "drawer_shell".
-The child link of this joint is "drawer_floor".
-It allows movement of the "drawer_floor" relative to the "drawer_shell," enabling the functionality of opening and closing the drawer.
-    "table":
-        type: actor
-        fixed: True
-        spatial_relationship: This is a fixed actor. The "drawer" and "apple" are on the "table".
-        description: None
-RESPONSE:
-stages:
-    stage_1:
-        name: open "drawer_floor" (left)
-        right_hand: empty
-        left_hand: empty
-        description: Open the drawer to access inside. This stage use left arm.
-    stage_2:
-        name: grasp "apple" (right)
-        right_hand: holding "apple"
-        left_hand: Grasp the apple on the table. This stage use right arm.
-    stage_3:
-        name: put "apple" into "drawer_floor" (right)
-        right_hand: empty
-        left_hand: empty
-        description: Put the apple into the drawer use right arm. This stage use right arm.
-Now, let’s begin:
-E.3 Baseline: Prompt for TP-D
-{mdframed}
-[backgroundcolor=azure, linewidth=0pt]
-You are a helpful assistant to plan a definite task into multi-satge for a dual-arm robot. To be on efficiency side, you can use dual arm each stage.
-I will give you the initial environment description:
-task: …
-obj_names: […]
-obj_description: …
-You must follow the following criteria:
-1) Break down the task into stages. The task can be complex but should be broken into simple, reasonable stages.
-2) The stage name should adhere the format: "API (right) and API (left)".
-The API must select in command API and follow the description.
-Note that right API and left API is executing at the same time, without order.
-3) If you need to use both hand too manipulate same object, the stage name should adhere the format: "API (both)".
-4) When moving an object, confirm it is grasped by hand and clearly specify its destination, e.g., ’put "obj_name" into/onto "obj_name"’.
-5) Synethesis current hand empty or holding obj after stage has been executed. For example, right_hand: holding "apple", left_hand: empty.
-6) The next stage you generate should accroding to hand.
-For example, right_hand: holding "apple", left_hand: empty.
-The stage_next: put "apple" into "drawer", is reasonable then your right hand is empty, you can use your right_hand.
-The stage_next: grasp "pear" is wrong because your right hand state is holding "apple".
-7) Each grasp stage should have a corresponding put stage which means a stage: grasp obj must have a corresponding stage: put obj.
-8) When the obj in your hand has been used up, it should be returned to a fitting position to release your hands.
-9) When two hands are needed to grasp or put an object, use a two-handed stage directly instead of two one-handed stages.
-For example, use stage: grasp pot (both) instead stage: grasp pot_left_handle (left) and grasp pot_right_handle (right).
-10) Enclose all noun references in quotes "". These must be in obj_names or ee_names. Do not introduce new objects or remove any.
-11) You need to generate as few stages as possible in order to complete the task as quickly as possible while still being able to reach the end goal.
-Any stage that is not relevant to the completion of the task should not be present.
-12) Please try to take advantage of the dual-arm robot’s two-armed feature.
-List of commands of the API, all the object must selected in obj_names:
-* The following command allows robot to open an <target-object> with a door, a drawer floor, a faucet:
-  command syntax:
-  -open <target-object>
-  -close <target-object>
-* The following command allows robot to switch on or off an electric device.
-If the electric device has a door, such as a mircrowave with door, this command can only be selected when the door of electric device is closed:
-  command syntax:
-  - switch_on <target-object>
-  - switch_off <target-object>
-* The following command allows robot to put an <object> in or on a <target-object>, the <object> must holding in hand, the <target-object> should not holding in hand:
-  command syntax:
-  - put <object> into <target-object>
-  - put <object> onto <target-object>
-{mdframed}[backgroundcolor=azure, linewidth=0pt]
-* The following command allows robot to pour a liquid from an <object> into an <target-object>. This command must be used to pour bottle of liquid:
-  command syntax:
-  - pour
-<object> into <target_object>
-* The following command allows robot to cut a food which is <target-object> with an <object>.
-This <target-object> must on the cutting_board, after cutting the <target-object> keep same name before cutting, and still on cutting board not in hand:
-  command syntax:
-  -cut <target-object> with <object>
-* The following command allows robot to do nothing:
-  command syntax:
-  - None
-You should only respond in the format as described below and do not respond anything else:
-RESPONSE FORMAT:
-stages:
-[…]
-        name: The name of the stage.
-        right_hand_state: Empty or holding sth.
-        left_hand_state: Empty or holding sth.
-Here’s an example input and response:
-INPUT:
-task: put "apple" into "drawer_floor"
-obj_names:  ["apple", "drawer_shell", "drawer_floor", "table"]
-obj_description:
-    "apple":
-        type: actor
-        fixed: False
-        spatial_relationship: The "apple" is on the "table". The "apple" is close to "right_hand".
-        description: None
-    "drawer":
-        type: articulation
-        fixed: True
-        links: ["drawer_shell","drawer_floor","drawer_handle"]
-        joints:  ["drawer_joint"]
-     spatial_relationship: This is a fixed articulation. The "drawer" is on the "table". The "drawer" is closed now. The "drawer" is close to "left_hand".
-        description:
-        "drawer_shell": This is the external part of the drawer, which cannot be manipulated but can be used to place items.
-The drawer shell is typically the external frame or casing of the drawer, providing structural support.
-        "drawer_floor": This is the internal part of the drawer, containing a "drawer_handle" and "drawer_joint." The drawer floor can be opened or closed by manipulating the "drawer_handle."
-         "drawer_handle": This is a handle located inside the drawer, and by operating it, the drawer can be opened or closed.
-Typically, the drawer handle is a component connected to the internal mechanism of the drawer.
-         "drawer_joint": The parent link of this joint is "drawer_shell".
-The child link of this joint is "drawer_floor".
-It allows movement of the "drawer_floor" relative to the "drawer_shell," enabling the functionality of opening and closing the drawer.
-    "table":
-        type: actor
-        fixed: True
-{mdframed}[backgroundcolor=azure, linewidth=0pt]         spatial_relationship: This is a fixed actor. The "drawer" and "apple" are on the "table".
-        description: None
-RESPONSE:
-stages:
-    stage_1:
-        name: grasp "apple" (right) and open "drawer_floor" (left)
-        right_hand: holding "apple"
-        left_hand: empty
-       description: Grasp the apple on the table use right arm and open the drawer to access inside use left arm.
-    stage_2:
-        name: put "apple" into "drawer_floor" (right) and None (left)
-        right_hand: empty
-        left_hand: empty
-        description: Put the apple into the drawer use right arm.
-E.4 Insturction and Environment Description of Task5
-{mdframed}
-[backgroundcolor=azure, linewidth=0pt] task: heat the soup and put the tin on the table
-constraints:
-    1) do not need to take bowl out of microwave after heating
-    2) close the door of refrigerator cooler after using
-obj_names:["table", "dishwasher","dishwasher_rack", "dishwasher_door", "microwave","microwave_body", "microwave_door", "tin", "bowl", "refrigerator_cooler_body", "refrigerator_cooler_door"]
-obj_description:
-    "table":
-        type: actor
-        fixed: True
-        spatial_relationship: The "table" is on the ground.
-        description: None
-    "dishwasher":
-        type: articulation
-        fixed: True
-        links: ["dishwasher_shell", "dishwasher_rack", "dishwasher_door", "dishwasher_handle", "dishwasher_switch"
-        joints: ["door_joint", "switch_joint"]
-       spatial_relationship: The "dishwasher" is under the "table". The "dishwasher" is at the left part of the "table".
-        description:
-            "dishwasher_shell": The shell of the "dishwasher".
-            "dishwasher_rack": The rack of the "dishwasher".
-            "dishwasher_door": The door of the "dishwasher".
-            "microwave_handle": The handle of the "dishwasher".
-            "dishwasher_switch": The switch of the "dishwasher".
-            "door_joint": The "door_joint" enables opening and closing of "dishwasher_door" by manipulating "microwave_handle".
-            "switch_joint": The "switch_joint" enables and disables of "dishwasher" by switching off and switching on "dishwasher_switch".
-The "dishwasher_switch" can only be manipulated when "dishwasher_door" is closing.
-    "microwave":
-        type: articulation
-        fixed: True
-{mdframed}[backgroundcolor=azure, linewidth=0pt]         links: ["microwave_body", "microwave_door","microwave_handle","microwave_switch"]
-       joints: ["door_joint","switch_joint"]
-      spatial_relationship: The "microwave" is on the "table". The "microwave" is at the right part of the "table". The "microwave" at the left side of the "refrigerator".
-        description:
-            "microwave_body": The body of the "microwave".
-            "microwave_door": The door of the "microwave".
-            "microwave_handle": The handle of the "microwave".
-            "microwave_switch": The switch of the "microwave".
-            "door_joint": The "door_joint" enables opening and closing of "microwave_door" by manipulating "microwave_handle".
-            "switch_joint": The "switch_joint" enables and disables of "microwave" by switching off and switching on "microwave_switch".
-The "microwave_switch" can only be manipulated when "microwave_door" is closing.
-    "bowl":
-        type: actor
-        fixed: False
-        spatial_relationship: The "bowl" is in the "refrigerator_cooler". The "bowl" is on the right side of the "tin".
-        description: The "bowl" is filled with soup.
-    "tin":
-        type: actor
-        fixed: False
-        spatial_relationship: The "tin" is in the "refrigerator_cooler". The "tin" is on the left side of the "bowl".
-        description: None
-    "refrigerator_cooler":
-        type: articulation
-        fixed: True
-   links: ["refrigerator_cooler_body", "refrigerator_cooler_door", "refrigerator_cooler_handle"]
-        joints: ["door_joint"]
-        spatial_relationship: The "refrigerator_cooler" is on the ground which at the right side the "table".
-The "refrigerator_cooler" at the right side of the "microwave".
-The "tin" and "bowl" both in the "refrigerator_cooler".
-        description:
-        "refrigerator_cooler_body": The body of the cooler floor of "refrigerator_cooler".
-"refrigerator_cooler_door": The door of the cooler floor of "refrigerator_cooler".
-"refrigerator_cooler_handle": The handle of the cooler floor door of "refrigerator_cooler".
-        "door_joint": The "door_joint" enables opening and closing of "refrigerator_cooler_door" by manipulating "refrigerator_cooler_handle".
-E.5 Insturction and Environment Description of Task8
-{mdframed}
-[backgroundcolor=azure, linewidth=0pt] task: cut carrot and red bell pepper
-constraints:
-    1) After cutting carrot and red bell pepper, put the knife onto the cutting board
-obj_names:["table", "sink", "faucet", "stove", "stove_body", "cutting_board", "cabinet_body", "cabinet_left_door", "cabinet_right_door", "knife_rack", "shelf_bottom_floor", "shelf_top_floor", "knife", "bowl", "pot", "carrot", "red_bell_pepper"]
-obj_description:
-    "table":
-        type: actor
-        fixed: True
-        spatial_relationship: The "table" is on the ground.
-        description: None
-    "sink":
-{mdframed}[backgroundcolor=azure, linewidth=0pt]         type: actor
-        fixed: True
-        spatial_relationship: The "sink" is on right side of the table.
-        description: A sink on the right side of the table. It offers a deep basin, can put the pot into it for filling water it by faucet.
-    "faucet":
-        type: articulation
-        fixed: True
-        links: ["faucet_body", "faucet_handle"]
-        joints: ["faucet_joint"]
-        spatial_relationship: The "faucet" is on the "table". The "faucet" is at the back of the sink.
-        description:
-        "faucet_body": "When pot in the sink and faucet is openning, pot can be filled with water."
-        "faucet_handle": "The faucet handle of faucet."
-        "faucet_joint": "The "faucet_joint" enables opening and closing of "faucet" by oepning and closing "faucet_handle". It provides flexibility in directing water flow."
-    "stove":
-        type: articulation
-        fixed: True
-        links: ["stove_body", "fire_knob"]
-        joints: ["stove_joint"]
-        spatial_relationship: The "stove" is on the "table". The "stove" is at the left side of "sink".
-        description:
-        "stove_body": "It allows which put on it for versatile cooking options."
-        "fire_knob": "A knob in front of the stove. The fire knob provides easy ignition control."
-        "stove_joint": "The "stove_joint" enables switching on and switching off of "stove" by manipulate "fire_knob"."
-    "cutting_board":
-        type: actor
-        fixed: True
-        spatial_relationship: The "cutting_board" is on the table.
-        description: "A cutting board on the left front side of the table. You can’t directly cut food in your hand. Any food must be put on the cutting board first before cutting."
-    "cabinet":
-        type: articulation
-        fixed: True
-        links: ["cabinet_body", "cabinet_door_left", "cabinet_left_handle", "cabinet_door_right", "cabinet_right_handle", "knife_rack"]
-        joints: ["cabinet_right_joint", "cabinet_left_joint"]
-      spatial_relationship: The "tool_cabinet" is above "table". The "cabinet_door_left" and "cabinet_door_right" is close.
-        description:
-         "cabinet_body": "It provides three storage space for bowl, knives and pot from left to right.
-The "bowl", "knife", and "knife_rack" is in the left part of "cabinet".
-The "pot" is in the right part of "tool_cabinet"."
-         "cabinet_left_door": "The left door of the cabinet, only it opened can access to "knife_rack"."
-        "cabinet_left_handle": "The left door is equipped handle for convenient opening and closing."
-        "cabinet_right_door": "The right door of the cabinet, only it opened can access to the right half of the cabinet."
-        "cabinet_right_handle": "The right door is equipped handle for convenient opening and closing."
-         "knife_rack": "The knife rack inside the left part of cabinet. It can securely hold kitchen knives. It has three slots for knives."
-{mdframed}[backgroundcolor=azure, linewidth=0pt]         "cabinet_right_joint": "The "cabinet_right_joint" enables opening and closing of "cabinet_door_right" by manipulating "cabinet_right_handle"."
-        "cabinet_left_joint": "The "cabinet_left_joint" enables opening and closing of "cabinet_door_left" by manipulating "cabinet_left_handle"."
-    "shelf":
-        type: articulation
-        fixed: True
-        links: ["shelf_bottom_floor", "shelf_top_floor"]
-        joints: None
-      spatial_relationship: The "shelf" is on the table. The "shelf" at the back side of "cutting_board".
-        description:
-            "shelf_bottom_floor": "The shelf has two floor. The "shelf_bottom" is the bottom floor of "shelf"."
-            "shelf_top_floor": "The "shelf_top" is the top floor of "shelf"."
-    "knife":
-        type: actor
-        fixed: False
-        spatial_relationship: "knife" is in "knife_rack".
-        description: None
-    "bowl":
-        type: actor
-        fixed: False
-        spatial_relationship: "bowl" is in "cabinet_body". "bowl" is at the left of "knife_rack".
-        description: None
-    "pot":
-        type: articulation
-        fixed: False
-        links: ["pot_body", "pot_left_handle", "pot_right_handle"]
-        joints: None
-        spatial_relationship: "pot" is in "cabinet_body".
-        description:
-            "pot_body": "The body of the "pot". It has a lot of mass and if you want to move it, you need to hold the handles with both hand."
-            "pot_left_handle": The left handle of the "pot".
-            "pot_right_handle": The right handle of the "pot".
-    "carrot":
-        type: actor
-        fixed: False
-        spatial_relationship: "carrot" is on "shelf_bottom_floor".
-        description: None
-    "red_bell_pepper":
-        type: actor
-        fixed: False
-        spatial_relationship: "red_bell_pepper" is on "shelf_bottom_floor".
-        description: None
-Appendix F Example Outputs
-F.1 Our method: DAG-Plan Output of Task 5
-{mdframed}
-[backgroundcolor=azure, linewidth=0pt] nodes:
-node_1:
-name: open "refrigerator_cooler_door"
-arm_num: 1
-edge: []
-type: operate
-description: Open the refrigerator cooler door to access the contents inside.
-node_2:
-name: grasp "bowl"
-arm_num: 1
-edge: [1]
-type: occupy
-description: Grasp the bowl containing soup from the refrigerator cooler.
-node_3: name: open "microwave_door"
-arm_num: 1
-edge: []
-type: operate
-description: Open the microwave door to place the bowl inside.
-node_4: name: put "bowl" into "microwave"
-arm_num: 1
-edge: [2, 3]
-type: release
-description: Place the bowl inside the microwave for heating.
-node_5: name: close "microwave_door"
-arm_num: 1
-edge: [4]
-type: operate
-description: Close the microwave door to start the heating process.
-node_6:
-name: switch_on "microwave"
-arm_num: 1
-edge: [5]
-type: operate
-description: Switch on the microwave to heat the soup in the bowl.
-node_7:
-name: switch_off "microwave"
-arm_num: 1
-edge: [6]
-type: operate
-description: Switch off the microwave after the soup is heated.
-node_8:
-name: grasp "tin"
-arm_num: 1
-edge: [1]
-type: occupy
-description: Grasp the tin from the refrigerator cooler.
-node_9:
-name: put "tin" onto "table"
-arm_num: 1
-edge: [8]
-type: release
-description: Place the tin onto the table.
-{mdframed}[backgroundcolor=azure, linewidth=0pt] node_10:
-name: close "refrigerator_cooler_door"
-arm_num: 1
-edge: [2, 8]
-type: operate
-description: Close the refrigerator cooler door after taking out the bowl and tin.
-node_11:
-name: task complete
-arm_num: 0
-edge: [7, 9, 10]
-type: complete
-description: Task complete after heating the soup and placing the tin on the table, and closing the refrigerator cooler door.
-F.2 Baseline: TP-S Output of Task 5
-{mdframed}
-[backgroundcolor=azure, linewidth=0pt] stages:
-stage_1:
-name: open "refrigerator_cooler_door" (left)
-right_hand_state: empty
-left_hand_state: empty
-description: Open the refrigerator cooler door to access the items inside. This stage uses the left arm.
-stage_2:
-name: grasp "tin" (right)
-right_hand_state: holding "tin"
-left_hand_state: empty
-description: Grasp the tin from the refrigerator cooler. This stage uses the right arm.
-stage_3: name: put "tin" onto "table" (right)
-right_hand_state: empty
-left_hand_state: empty
-description: Place the tin onto the table. This stage uses the right arm.
-stage_4:
-name: grasp "bowl" (right)
-right_hand_state: holding "bowl"
-left_hand_state: empty
-description: Grasp the bowl filled with soup from the refrigerator cooler. This stage uses the right arm.
-stage_5:
-name: open "microwave_door" (left)
-right_hand_state: holding "bowl"
-left_hand_state: empty
-description: Open the microwave door to place the bowl inside. This stage uses the left arm.
-stage_6:
-name: put "bowl" into "microwave_body" (right)
-right_hand_state: empty
-left_hand_state: empty
-description: Place the bowl inside the microwave for heating. This stage uses the right arm.
-stage_7:
-name: close "microwave_door" (left)
-right_hand_state: empty
-left_hand_state: empty
-description: Close the microwave door after placing the bowl inside. This stage uses the left arm.
-{mdframed}[backgroundcolor=azure, linewidth=0pt] stage_8:
-name: switch_on "microwave" (right)
-right_hand_state: empty
-left_hand_state: empty
-description: Switch on the microwave to start heating the soup. This stage uses the right arm.
-stage_9:
-name: switch_off "microwave" (right)
-right_hand_state: empty
-left_hand_state: empty
-description: Switch off the microwave after the soup is heated. This ensures safety by keeping the appliance off after use. This stage uses the right arm.
-stage_10:
-name: close "refrigerator_cooler_door" (left)
-right_hand_state: empty
-left_hand_state: empty
-description: Close the refrigerator cooler door after retrieving and returning items. This stage uses the left arm.
-F.3 Baseline: TP-D Output of Task 5
-{mdframed}
-[backgroundcolor=azure, linewidth=0pt]
-stages:
-stage_1:
-name: open "refrigerator_cooler_door" (left) and open "microwave_door" (right)
-right_hand_state: empty
-left_hand_state: empty
-description: Open the refrigerator cooler door with left hand and microwave door with right hand to access the bowl and microwave.
-stage_2:
-name: grasp "bowl" (left) and grasp "tin" (right)
-right_hand_state: holding "tin"
-left_hand_state: holding "bowl"
-description: Grasp the bowl with soup from the refrigerator with left hand and the tin with right hand.
-stage_3:
-name: put "bowl" into "microwave_body" (left) and put "tin" onto "table" (right)
-right_hand_state: empty
-left_hand_state: holding "bowl"
-description: Place the bowl inside the microwave with left hand and put the tin on the table with right hand.
-stage_4:
-name: close "refrigerator_cooler_door" (right) and switch_on "microwave" (left)
-right_hand_state: empty
-left_hand_state: empty
-description: Close the refrigerator cooler door using the right hand and switch on the microwave with the left hand to start heating the soup.
-stage_5:
-name: switch_off "microwave" (left) and None (right)
-right_hand_state: empty
-left_hand_state: empty
-description: Turn off the microwave after heating the soup to ensure safety, with the left hand.
-F.4 Our method: DAG-Plan Output of Task 8
-{mdframed}
-[backgroundcolor=azure, linewidth=0pt] nodes:
-node_1:
-name: open "cabinet_left_door"
-arm_num: 1
-edge: []
-type: operate
-description: Open the left door of the cabinet to access the knife rack.
-node_2:
-name: grasp "knife"
-arm_num: 1
-edge: [1]
-type: occupy
-description: Grasp the knife from the knife rack.
-node_3:
-name: grasp "carrot"
-arm_num: 1
-edge: []
-type: occupy
-description: Grasp the carrot from the shelf bottom floor.
-node_4:
-name: put "carrot" onto "cutting_board" arm_num: 1
-edge: [3]
-type: release
-description: Place the carrot onto the cutting board to prepare for cutting.
-node_5:
-name: cut "carrot" with "knife"
-arm_num: 1
-edge: [2, 4]
-type: tool use
-description: Cut the carrot using the knife while it is on the cutting board.
-node_6:
-name: grasp "red_bell_pepper"
-arm_num: 1
-edge: []
-type: occupy
-description: Grasp the red bell pepper from the shelf bottom floor.
-node_7:
-name: put "red_bell_pepper" onto "cutting_board"
-arm_num: 1
-edge: [6]
-type: release
-description: Place the red bell pepper onto the cutting board to prepare for cutting.
-node_8:
-name: cut "red_bell_pepper" with "knife"
-arm_num: 1
-edge: [2, 7]
-type: tool use
-description: Cut the red bell pepper using the knife while it is on the cutting board.
-node_9:
-name: put "knife" onto "cutting_board"
-arm_num: 1
-edge: [5, 8]
-type: release
-description: Place the knife back onto the cutting board after use.
-{mdframed}[backgroundcolor=azure, linewidth=0pt] node_10:
-name: task complete
-arm_num: 0
-edge: [9]
-type: complete
-description: Task complete after all cutting and placements are done.
-F.5 Baseline: TP-S Output of Task 8
-{mdframed}
-[backgroundcolor=azure, linewidth=0pt] stages:
-stage_1:
-name: open "cabinet_left_door" (left)
-right_hand_state: empty
-left_hand_state: empty
-description: Open the left door of the cabinet to access the knife rack.
-stage_2:
-name: grasp "knife" (right)
-right_hand_state: holding "knife"
-left_hand_name: empty
-description: Grasp the knife from the knife rack.
-stage_3:
-name: close "cabinet_left_door" (left)
-right_hand_state: holding "knife"
-left_hand_state: empty
-description: Close the left door of the cabinet after taking out the knife.
-stage_4:
-name: grasp "carrot" (left)
-right_hand_state: holding "knife"
-left_hand_state: holding "carrot"
-description: Grasp the carrot from the shelf bottom floor.
-stage_5:
-name: put "carrot" onto "cutting_board" (left)
-right_hand_state: holding "knife"
-left_hand_state: empty
-description: Place the carrot onto the cutting board to prepare for cutting.
-stage_6:
-name: cut "carrot" with "knife"
-right_hand_state: holding "knife"
-left_hand_state: empty
-description: Cut the carrot on the cutting board using the knife.
-stage_7:
-name: grasp "red_bell_pepper" (left)
-right_hand_state: holding "knife"
-left_hand_state: holding "red_bell_pepper"
-description: Grasp the red bell pepper from the shelf bottom floor.
-stage_8:
-name: put "red_bell_pepper" onto "cutting_board" (left)
-right_hand_state: holding "knife"
-left_hand_state: empty
-description: Place the red bell pepper onto the cutting board to prepare for cutting.
-stage_9:
-name: cut "red_bell_pepper" with "knife"
-right_hand_state: holding "knife"
-left_hand_state: empty
-description: Cut the red bell pepper on the cutting board using the knife.
-{mdframed}[backgroundcolor=azure, linewidth=0pt] stage_10:
-name: put "knife" onto "cutting_board" (right)
-right_hand_state: empty
-left_hand_state: empty
-description: Place the knife back onto the cutting board after cutting the vegetables.
-F.6 Baseline: TP-D Output of Task 8
-{mdframed}
-[backgroundcolor=azure, linewidth=0pt]
-stages:
-stage_1:
-name: grasp "carrot" (right) and grasp "red_bell_pepper" (left)
-right_hand_state: holding "carrot"
-left_hand_state: holding "red_bell_pepper"
-description: The robot grasps the "carrot" with its right hand and the "red_bell_pepper" with its left hand from the "shelf_bottom_floor".
-stage_2:
-name: put "carrot" onto "cutting_board" (right) and put "red_bell_pepper" onto "cutting_board" (left)
-right_hand_state: empty
-left_hand_state: empty
-description: The robot places both the "carrot" and "red_bell_pepper" onto the "cutting_board" to prepare them for cutting.
-stage_3:
-name: open "cabinet_left_door" (left) and None (right)
-right_hand_state: empty
-left_hand_state: empty
-description: The robot opens the left door of the "cabinet" to access the "knife_rack".
-stage_4:
-name: grasp "knife" (right) and None (left)
-right_hand_state: holding "knife"
-left_hand_state: empty
-description: The robot grasps the "knife" from the "knife_rack".
-stage_5:
-name: cut "carrot" with "knife" (right) and cut "red_bell_pepper" with "knife" (left)
-right_hand_state: holding "knife"
-left_hand_state: empty
-description: The robot uses the "knife" to cut both the "carrot" and the "red_bell_pepper" on the "cutting_board".
-stage_6:
-name: put "knife" onto "cutting_board" (right) and None (left)
-right_hand_state: empty
-left_hand_state: empty
-description: After cutting, the robot places the "knife" back onto the "cutting_board" as per the task constraints.
+The main contribution is the replacement of task sequence with a DAG and the dynamic adjustment of the planning according to the current situation, allowing dual-arm robots to flexibly utilize both arms for sub-task execution.
+Extensive experiments demonstrate that DAG-Plan integrates the advantages of various methods, exhibiting balanced and superior performance.
+Compared to the single-arm approach, DAG-Plan demonstrates significantly higher execution efficiency.
+In contrast to dual-arm methods, DAG-Plan achieves a higher success rate in task planning.
+When compared to the iterative method, DAG-Plan maintains the ability to interact with environmental information while requiring shorter query time and reduced token usage.
+ACKNOWLEDGEMENTS
+The authors would like to thank Qiaojun Yu, Tianxing Chen for their fruitful discussions throughout the project and for providing helpful feedback on initial drafts of the manuscript.